### v2

# Yuanyuan Wang1    Tianze Wei2

Abstract
In an online fair allocation problem, a sequence of indivisible items arrives online and needs to be allocated to offline agents immediately and irrevocably. In our paper, we study the online allocation of either goods or chores. We employ popular fairness notions, including envy-freeness up to one item (EF1) and maximin share fairness (MMS) to capture fairness, and utilitarian social welfare (USW) to measure efficiency. For both settings of items, we present a series of positive results regarding the existence of fair and efficient allocations with widely studied classes of additive binary and personalized bi-valued valuation/cost functions. Furthermore, we complement our results by constructing counterexamples to establish our results as among the best guarantees possible.
1 Introduction
The fair division of indivisible items is a prominent topic in algorithmic game theory and artificial intelligence, with practical applications Budish (2011); Goldman and Procaccia (2015). Most prior work has focused on offline settings, where the information of all items is known a priori. In practice, items are often online, and waiting for all resources to appear can significantly impede progress. We must proactively allocate resources in a timely manner for better outcomes. A paradigmatic example is the food banks problem Lee et al. (2019), where food arrives online without prior knowledge of future items. Since the food is perishable, it must be allocated upon arrival to some agent. Motivated by these real-world scenarios, we define and discuss the following problem:
(Online Fair Division) There are $n$ agents that are fixed and known in advance. The items arrive online (the number of items is unknown). Upon the arrival of an item, its values to all agents are revealed, and we must allocate it to some agent or discard it immediately and irrevocably. At all times, the allocation among agents must be fair.
If the online algorithm knows additional information about all future items beforehand, it somewhat reduces the uncertainty of the online model, allowing for more effective decision-making and improved outcomes. Zhou et al. (2023); Benadè et al. (2024) study the normalized valuations, i.e., online algorithms know the total value of future items. Benadè et al. (2022) studies the online fair allocation, foreseeing the rank of a new arrival item with respect to the past items under different assumptions of agents’ valuation distributions. And there are other studies about additional information He et al. (2019); Elkind et al. (2024); Cookson et al. (2024). For an independent interest, we also examine the online setting with additional information, with relevant results presented in the Appendix.
In numerous practical applications, such as vaccine allocation Rey et al. (2023), the online algorithm operates without prior knowledge of future items. Aleksandrov et al. (2015) studies the online fair division with additive binary valuations and shows that an envy-free up to one item (EF1) allocation exists. However, there are a few non-trivial approximation fairness results of online goods allocation beyond additive binary valuations and online chores allocation. In our paper, we aim to address this gap and study the online allocation model without any information about future items. We discuss the goods and chores settings, respectively. Our primary research question is:
Can we design deterministic online algorithms that are fair and efficient, for valuation functions beyond additive binary, in the online allocation problem?
1.1 Our Contributions
We study two settings: goods and chores, in the online fair indivisible items allocation model.
In the goods setting, for different classes of valuations, including binary and additive personalized bi-valued, our main results are summarized in Table 1. We introduce the non-wastefulness (NW) notion in this setting. First, we show the inapproximability of EF1 and MMS with general additive valuations, even for only two agents with additive personal tri-valued valuations. Then, we study some restricted valuation functions, such as submodular binary valuation functions.
Valuations
Number
of agents
Fairness/
Efficiency
LB UB
Additive
(personalized tri-valued)
[…]
Submodular
binary
$n$ EF1
$\frac{1}{2}$
(Thm 2)
$\frac{1}{2}$
(Thm 3)
[…]
MMS
$\frac{1}{2}$
(Thm 2)
$\frac{1}{2}$
(Thm 3)
[…]
USW
$\frac{1}{2}$
(Thm 2)
$\frac{1}{2}$
(Thm 3)
[…]
Additive personalized
bi-valued
22 EF1
$\frac{1}{2}$
(Thm 4)
$\frac{1}{2}$
(Thm 5)
[…]
MMS
$\frac{1}{3}$
(Thm 4)
$\frac{1}{3}$
(Thm 5)
[…]
Identical additive
binary
+ personalized bi-valued
[…]
USW $\times$ (Prop 3)
Table 1: Goods setting. LB: lower bound. UB: upper bound. $\times"$ indicates inapproximability.
  • For submodular binary valuation functions, we propose the Marginal-Greedy Algorithm (Algorithm 2) to compute an allocation that simultaneously satisfies NW, $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW, and we show that these approximation ratios are all tight.
  • For two agents with additive personalized bi-valued valuation functions, we design the Adapted Envy-Graph Procedure (Algorithm 3) to compute an NW allocation satisfying $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS simultaneously. Further, we show that no deterministic algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\beta$-MMS for any $\alpha>\frac{1}{2}$, $\beta>\frac{1}{3}$, and no deterministic algorithm can compute an NW allocation where $\alpha$-EF1 (or $\alpha$-MMS) is compatible with $\beta$-max-USW for any $\alpha,\beta>0$.
  • When one agent has the additive personalized bi-valued valuation function and the remaining agents have identical additive binary valuation functions, we design the Adapted-Picking Algorithm (Algorithm 7) to compute an allocation satisfying NW, EF1, and MMS. Further, we show that no deterministic algorithm can compute an NW allocation that is EF1 (or MMS) and $\alpha$-max-USW for any $\alpha>0$.
Costs
Number
of agents
Fairness/
Efficiency
LB UB
Additive
(personalized tri-valued)
[…]
Additive
binary
[…]
Supermodular
binary
[…]
Additive personalized
bi-valued
22 EF1
22
(Thm 11)
22
(Thm 10)
[…]
MMS
$\frac{3}{2}$
(Thm 11)
$\frac{5}{3}$
(Thm 10)
[…]
Additive binary
+ personalized bi-valued
[…]
Table 2: Chores setting. LB: lower bound. UB: upper bound. $\times"$ indicates inapproximability.
In the chores setting, we introduce the completeness constraints. We first show the inapproximability of EF1 even for two agents with additive personalized tri-valued cost functions, which blocks our way of studying more general cost functions. Next, we investigate the restricted cost functions, such as additive binary cost functions.
  • For additive binary cost functions, we propose the Compelled-Greedy Algorithm (Algorithm 8) to compute a complete allocation that simultaneously satisfies EF1, MMS, and min-USC. For supermodular binary cost functions, we prove that no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 (or $\alpha$-MMS or $\alpha$-min-USC) for any $\alpha\geq 1$.
  • When two agents have additive personalized bi-valued cost functions, we design the Adapted Chores Envy-Graph Procedure (Algorithm 9) to compute an allocation that satisfies completeness, $2$-EF1, and $\frac{5}{3}$-MMS. Further, we show that no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any $\alpha<2$ or $\beta$-MMS for any $\beta<\frac{3}{2}$, and no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 (or $\beta$-MMS) and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$.
  • When one agent has the additive personalized bi-valued cost function and the remaining agents have additive binary cost functions, we design the Adapted-Chores-Picking Algorithm (Algorithm 13) to compute an allocation satisfying completeness, EF1, and MMS. Further, no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 (or MMS) and $\beta$-min-USC for any $\alpha,\beta\geq 1$.
1.2 Related Work
Online Fair Allocation.
In the majority of the literature, there are two kinds of online arrivals: The former kind is that agents arrive over time Kash et al. (2014); Friedman et al. (2015, 2017); Li et al. (2018), and the latter is that items arrive over time, like Aleksandrov et al. (2015), which is also the topic of our paper. Our work is mainly related to the study of Aleksandrov et al. (2015) and Hosseini et al. (2024), where items arrive one by one without any future information known in advance. Aleksandrov et al. (2015) consider the food bank problem when all agents only have the additive binary valuation functions and design two simple mechanisms to guarantee ex-ante envy-freeness. In our paper, we study beyond additive binary valuation functions in the goods setting. Hosseini et al. (2024) study the class fairness in online matching for indivisible and divisible items, and present a series of positive results. The model of indivisible items in their paper can be seen as a special case of our model with submodular binary valuation functions in our paper, which will be discussed later. In other words, we study a more general setting and derive the approximation results that strengthen their results.
Offline Fair Allocation.
When agents have additive valuation functions, there is a rich body of literature that discusses EF1 or MMS. Please refer to the survey by Amanatidis et al. (2023) for an overview. Here, we highlight some papers most relevant to our work on offline fair division with submodular binary valuation functions or supermodular binary cost functions. When agents have submodular binary valuations in the goods setting, Benabbou et al. (2021) study the compatibility of fairness and efficiency and show that a maximum utilitarian social welfare allocation that is EF1 can be efficiently computed, and Barman and Verma (2021) consider the fairness notion - MMS and prove that a maximum utilitarian social welfare allocation that satisfies MMS can be computed in polynomial time. When agents have supermodular binary cost functions in the chores setting, Barman et al. (2023) show that an EF1 or MMS allocation minimizing the social cost can be computed in polynomial time.
2 Preliminaries
For each natural number $s\in\mathbb{N}$, let $s]=\{1,\ldots,s\}$. Let $T=\{e_{1},\ldots,e_{t}\}$ be the set of indivisible items arriving one by one and $N=\{1,\ldots,n\}$ be the set of offline agents. When an online item $e_{t}$ arrives, its values or costs to all agents are revealed. When items are goods, each agent is endowed with a valuation function $v_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. When items are chores, each agent is endowed with a cost function $c_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. We assume that $v_{i}(\emptyset)=0$ ($c_{i}(\emptyset)=0$), and monotone, $v_{i}(X)\leq v_{i}(Y)$ ($c_{i}(X)\leq c_{i}(Y)$) for all $X\subseteq Y\subseteq T$. For any item $e\in T$ and any subset $X\subseteq T$, we denote by $\Delta^{i}_{X}(e)=v_{i}(X\cup\{e\})-v_{i}(X)$ or $c_{i}(X\cup\{e\})-c_{i}(X)$ the marginal value or cost of an item $e\in T$ in the set $X$. $v_{i}$ is additive if the value of any set of items $S$ is the sum of the value of each item in $S$, that is, $v_{i}(X)=\sum_{e\in X}v_{i}(e)$. $v_{i}$ is submodular if $v_{i}(X)+v_{i}(Y)\geq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. $v_{i}$ is supermodular if $v_{i}(X)+v_{i}(Y)\leq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. There are similar arguments for $c_{i}$, and we omit it. Let $T^{k}$ be the set of items that have arrived by round $k\in[t]$. We denote an allocation by ${\bf A}^{k}=(A_{1}^{k},\ldots,A_{n}^{k})$, where $A_{i}^{k}\subseteq T^{k}$ is the bundle of items that are allocated to agent $i$, and $A_{i}^{k}\cap A_{j}^{k}=\emptyset$ for any $i,j\in N$. Let $\Pi_{n}(T^{k})$ be the set of all allocations of $T^{k}$. Collectively, $(T,N,(v_{i})_{i\in N})$ forms an instance of online goods fair allocation and $(T,N,(c_{i})_{i\in N})$ forms an instance of online chores fair allocation. For a singleton set $\{e\}$, we will use $v_{i}(e)$ ($c_{i}(e)$) as a shorthand for $v_{i}(\{e\})$ ($c_{i}(\{e\})$).
Deterministic online algorithms.
In the online setting, the items in $T$ arrive one by one in an arbitrary order, where $|T|$ is unknown. When an item arrives, the values (costs) of this item for all agents are revealed. The online algorithm must make an immediate and irrevocable decision to allocate the item to one agent or discard it deterministically.
Our objective is to design deterministic algorithms to compute allocations guaranteeing the desired approximation fairness or efficiency, including EF1, MMS, and USW/USC (Definitions 1, 2, 3, 8, 9, and 10), as well as constraints, including non-wastefulness (Definition 4) and completeness (Definition 7), at the end of each round. In contrast, our impossibility results will hold even if the desired guarantees are required to hold only at the end. For a better understanding of our online model, let us demonstrate the deterministic online allocation procedure through the following example.
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 6 88 10 66
agent 2 13 8 5 0
Table 3: Example of the deterministic online allocation.
Example.
Suppose that there are two agents who have additive valuation functions, and the arriving items are goods. The value of each item is shown in Table 3, and note that the value of each item is only revealed after it arrives. Now we have a deterministic online algorithm named $\mathsf{OALG}$. In round $t=1$, $\mathsf{OALG}$ allocates $e_{1}$ to agent 1, and the allocation ${\bf A}^{1}$ is EF1, MMS, and $\frac{6}{13}$-max-USW. In round $t=2$, $\mathsf{OALG}$ allocates $e_{2}$ to agent 2, and the allocation ${\bf A}^{2}$ is EF1, MMS, and $\frac{2}{3}$-max-USW. In round $t=3$, $\mathsf{OALG}$ allocate $e_{3}$ to agent 2, and the allocation $\textbf{A}^{3}$ is $\frac{3}{4}$-EF1, $\frac{3}{5}$-MMS, and $\frac{19}{31}$-max-USW. In round $t=4$, $\mathsf{OALG}$ discards $e_{4}$, and the allocation $\textbf{A}^{4}$ is $\frac{3}{4}$-EF1, $\frac{3}{7}$-MMS, and $\frac{19}{37}$-max-USW. Then, no further item arrives and we say that $\mathsf{OALG}$ achieves $\min\{\frac{3}{4},1\}=\frac{3}{4}$-EF1, $\min\{\frac{3}{7},\frac{3}{5},1\}=\frac{3}{7}$-MMS, and $\min\{\frac{6}{13},\frac{19}{37},\frac{19}{31},\frac{2}{3}\}$$\frac{6}{13}-$max-USW.
3 Allocation of Goods
Throughout this section, we focus on the goods setting. We first introduce some definitions of fairness and efficiency notions that are used in this section.
Definition 1 (Envy-freeness up to One Good).
For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate envy-free up to one good ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i^{\prime}}^{k}\neq\emptyset$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot v_{i}(A_{i^{\prime}}^{k}\setminus\{e\})$ for some $e\in A_{i^{\prime}}^{k}$.
Definition 2 (Maximin Share Fairness).
For any agent $i$, her maximin share $\mathsf{MMS}_{i}^{k}$ by round $t$ is defined as:
$\mathsf{MMS}_{i}^{k}=\max_{B^{k}\in\Pi_{n}(T^{k})}\min_{i^{\prime}\in N}v_{i}(B_{i^{\prime}}^{k}),$
For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate maximin share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
Definition 3 (Utilitarian Social Welfare).
The utilitarian social welfare of allocation ${\bf A}^{k}$ is given by $USW({\bf A}^{k})=\sum_{i\in N}v_{i}(A_{i}^{k})$. For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-max-USW if $USW({\bf A}^{k})\geq\alpha\cdot\max_{{\bf B}^{k}\in\Pi_{n}(T^{k})}USW({\bf B}^{k})$.
Besides the above definitions, we also pay attention to the following restriction, which is added in the goods setting.
Definition 4 (Non-Wastefulness).
An allocation ${\bf A}^{k}$ is non-wasteful (NW) if (1) $\Delta_{A_{i}^{k}\setminus\{e\}}^{i}(e)>0$ for any agent $i\in N$ and any good $e\in A^{k}_{i}$, and (2) $\Delta_{A_{i}^{k}}^{i}({e}^{{}^{\prime}})=0$ for any agent $i\in N$ and any discarded good $e^{\prime}\in T^{k}$.
Note that imposing NW in the goods setting is necessary, as our online model permits item discarding (see our example for illustration). Without this constraint, achieving EF1 would be trivial by simply discarding all online items. For agents with additive valuation functions, an NW allocation can always be found as long as each arriving item is allocated to at least one agent who values it positively. Below, we present a positive result where an NW allocation always exists for submodular binary valuation functions that we study in the subsequent parts, and complement it with a negative result for general submodular valuation functions.
Lemma 1.
For submodular binary valuation functions, an NW allocation always exists.
Proof.
Input: An instance $(T,N,(v_{i})_{i\in N})$ with submodular binary valuation functions
Output: A NW allocation $\bf A$}
1 Initialize: Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
2 when item $e\in T$ arrives do
3
4 for $i=1$ to $n$ do
5
6    if $\Delta_{A_{i}}^{i}({e})>0$ then
7
8        Let $A_{i}=A_{i}\cup\{e\}$;
9        break;
10      end if
11
12   end for
13
14
15
16return $\bf A$};
Algorithm 1 Greedy Algorithm
Given an instance $(T,N,(v_{i})_{i\in N})$, Algorithm 1 can compute an NW allocation. When item $e_{k}$ ($k\in[t]$) arrives, if there exists an agent $j\in N$ such that $\Delta_{A_{j}^{k-1}}^{j}({e_{k}})=1$, $e_{k}$ will be allocated. We assume that $e_{k}$ is allocated to agent $i$. We will show that $\Delta_{A_{i}^{r}\setminus\{e_{r}\}}^{i}({e_{k}})=1$ holds for any round $r>k$ after some new items are added to agent $i$’s bundle. Without loss of generality, we assume that $e_{s}$ is an arbitrarily newly added item in agent $i$’s bundle, where $s>k$, it holds that $\Delta_{A_{i}^{s-1}}^{i}({e_{s}})=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1})=1$. Then, we have
$\Delta_{{A_{i}^{s}}\setminus\{e_{r}\}}^{i}({e_{k}})$ $=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$=v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$\geq v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\setminus\{e_{k}\})-v_{i}(e_{s})$
$=v_{i}(A_{i}^{s-1})-v_{i}(A^{s-1}\setminus\{e_{k}\})$
$=\Delta_{A_{i}^{s-1}\setminus\{e_{k}\}}^{i}({e_{k}}),$
where the inequality holds for submodularity. Following this induction by repeatedly utilizing submodularity, we have that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})\geq\ldots\geq\Delta_{A_{i}^{k}\setminus\{e_{k}\}}^{i}({e_{k}})=\Delta_{A_{i}^{k-1}}^{i}({e_{k}})=1$. Since $v_{i}$ is a submodular binary valuation function, it holds that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})\leq 1$. Thus, we can conclude that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})=1$. ∎
Lemma 2.
For submodular valuation functions, there exists an instance that no deterministic online algorithm can compute an NW allocation, even for two agents with identical valuation functions.
Proof.
We construct an instance with two agents, $v_{1}=v_{2}=v$. When first item $e_{1}$ arrives, $v(e_{1})=3$. W.l.o.g., we allocate $e_{1}$ to agent 1. When the second item $e_{2}$ arrives, $v(e_{2})=3,v(\{e_{1},e_{2}\})=5$. Case (1): we allocate $e_{2}$ to agent 2, for the third item $e_{3}$$v(e_{3})=5,v(\{e_{1},e_{3}\})=v(\{e_{2},e_{3}\})=v(\{e_{1},e_{2},e_{3}\})=5$. We have that $v(\{e_{1},e_{3}\})-v(e_{3})=v(\{e_{2},e_{3}\})-v(e_{3})=0$. So, no matter who is assigned $e_{3}$, it cannot satisfy NW. Case (2): we allocate $e_{2}$ to agent 1, for the third item $e_{3}$ with $v(e_{3})=5,v(\{e_{1},e_{3}\})=v(\{e_{2},e_{3}\})=v(\{e_{1},e_{2},e_{3}\})=5$. To ensure NW, we must allocate $e_{3}$ to agent 2. For the fourth item $e_{4}$, $v(e_{4})=v(\{e_{1},e_{4}\})=v(\{e_{2},e_{4}\})=v(\{e_{3},e_{4}\})=v(\{e_{1},e_{2},e_{4}\})=v(\{e_{1},e_{3},e_{4}\})=v(\{e_{2},e_{3},e_{4}\})=v(\{e_{1},e_{2},e_{3},e_{4}\})=6$. We have that $v(\{e_{1},e_{2},e_{4}\})-v(\{e_{2},e_{4}\})=v(\{e_{3},e_{4}\})-v(e_{4})=0$. So, no matter who is assigned $e_{4}$, it cannot be NW. ∎
3.1 The Impossibilities of Approximate EF1 and MMS
(Aleksandrov et al., 2015) show that in the online setting, an EF1 allocation exists for additive binary valuations, where for each agent $i\in N$, $v_{i}$ is additive and $v_{i}(e)\in\{0,1\}$ for any $e\in T$. In this part, we present a strong negative result regarding the approximations of EF1 and MMS for general additive valuations. We say that the valuation function $v_{i}$ is additive personalized tri-valued if $v_{i}$ is additive and $v_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
Theorem 1.
No deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$, even for two agents with additive personalized tri-valued valuation functions.
Proof.
$e_{1}$ $e_{2}$ $e_{3}$
agent $1$ 1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
agent 2 11 $\epsilon$ $\frac{1}{\epsilon}$
Table 4: The impossibility result of EF1 and MMS.
Given any $\alpha>0$, let $\epsilon$ be a constant arbitrarily close to 0 such that $0<\epsilon<\alpha$. Consider an online instance with two agents. For any agent $i\in N$ and item $e\in T$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $0<\epsilon<1$. The items arrive online, and the valuation of each item is shown in Table 4. When $t=1$, the item $e_{1}$ arrives, $T^{1}=\{e_{1}\}$. Without loss of generality, suppose it is allocated to agent 1, $A^{1}_{1}=\{e_{1}\},A^{1}_{2}=\emptyset$. When $t=2$, the item $e_{2}$ arrives, $T^{2}=\{e_{1},e_{2}\}$. To maintain the fairness of allocation, we must allocate it to agent 2, i.e., $A^{2}_{1}=\{e_{1}\},A^{2}_{2}=\{e_{2}\}$. Otherwise, agent 1 gets $\{e_{1},e_{2}\}$ and agent 2 gets an empty set, $v_{2}(\emptyset)=0<v_{2}(\{e_{1},e_{2}\}\setminus\{e\})$ for any $e\in\{e_{1},e_{2}\}$ and $MMS_{2}(T^{2})=\epsilon$. It is 0-EF1 and 0-MMS. When $t=3$, item $e_{3}$ arrives, $T^{3}=\{e_{1},e_{2},e_{3}\}$. $MS_{1}(T^{3})=\frac{1}{\epsilon}, MM$$MS_{2}(T^{3})=1+\epsilon$. If we allocate it to agent 1, i.e., $A^{3}_{1}=\{e_{1},e_{3}\},A^{3}_{2}=\{e_{2}\}$, we have $v_{2}(A^{3}_{2})=\epsilon$, $v_{2}(A^{3}_{1}\setminus\{e_{1}\})=\frac{1}{\epsilon}$, and $v_{2}(A^{3}_{1}\setminus\{e_{3}\})=1$. Thus, $\bf A^{3}$} is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If $e_{3}$ is allocated to agent 2, i.e., $A^{3}_{1}=\{e_{1}\},A^{3}_{2}=\{e_{2},e_{3}\}$, we have $v_{1}(A^{3}_{1})=1=\epsilon\cdot v_{1}(A^{3}_{2}\setminus\{e\})$ for any $e\in A^{3}_{2}$. Thus, $\bf A^{3}$} is $\epsilon$-EF1 and $\epsilon$-MMS.
Therefore, in the above instance, no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS. This result can be generalized to more than two agents, and we defer the proof to Proposition 1. ∎
Proposition 1.
For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$ when there are more than two agents with additive personalized tri-valued valuation functions.
Proof.
$e_{1}$ $e_{2}$ $\ldots$ $e_{n-2}$ $e_{n-1}$ $e_{n}$ $e_{n+1}$
agent 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent 22 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent n2n-2 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent n1n-1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
Table 5: The impossibility results of EF1 and MMS for additive personalized tri-valued valuations when there are more than two agents.
We further extend the impossibility results of approximate EF1 and MMS for $n\geq 3$. Consider the following instance with $n\geq 3$ agents, where the value of each item is shown in Table 5. In the first (n2)(n-2) rounds, to avoid 0-EF1 or 0-MMS, each item is allocated to the agent with the empty bundle. For the item that arrives before round t=n2t=n-2, it has the value of $\frac{1}{\epsilon}$ for any agent. Without loss of generality, we assume that agent $i$ picks the item in the round $t=i$, where $i\leq n$. In round t=n1t=n-1, the arriving item has the value of 1 for any agent, and we assume that it is allocated to agent n1n-1. In round $t=n$, the arriving item has the value of $\epsilon$ for agent $i\in N\setminus\{n-1\}$ and $\frac{1}{\epsilon}$ for agent n1n-1, and it should be allocated to agent $n$ to avoid 0-EF1 or 0-MMS. Finally, in the round $t=n+1$, the item with the value of $\frac{1}{\epsilon}$ for each agent arrives. Due to non-wastefulness, this item must be allocated to one agent. If this item is allocated to agent $i\in N\setminus\{n-1,n\}$, it is not hard to see that the allocation is $\epsilon^{2}$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent n1n-1, this allocation is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent $n$, we can derive that the competitive ratio of EF1 is $\epsilon$ and MMS is $\epsilon$. Combining the above cases, we can conclude that the upper bound of EF1 is $\epsilon$ and MMS is $\epsilon$. For given $\alpha>0$, let $\epsilon<\alpha$, we can conclude that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$. ∎
By Theorem 1, we cannot make further progress in the approximations of EF1 and MMS for additive personalized tri-valued valuation functions and beyond in the online setting. In the following parts, we focus on two classes of valuation functions that are more general than additive binary functions, including submodular binary and additive personalized bi-valued valuation functions, and study whether there are positive results about the approximations of EF1 and MMS.
3.2 Submodular Binary Valuations
This part focuses on submodular binary valuation functions, where for each agent $i\in N$, $v_{i}$ is submodular and the marginal value of any item is 0 or 1. For a subclass of submodular binary valuation functions based on maximum (unweighted) bipartite matching, also known as assignment Munkres (1957) or OXS Lehmann et al. (2001) valuations, Hosseini et al. (2024) show that online allocations satisfying $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS exist.
Limitation of Hosseini et al. (2024): Their $\frac{1}{2}$-EF1/MMS guarantees are restricted to OXS valuations, where items must be assigned within predefined groups under rigid matching constraints. Each member in the group can only receive at most one item, and once someone gets an item, it cannot be changed. Thus, their algorithm can only handle the OXS valuation function and its subclasses.
Moving beyond the online allocation model based on matching, we consider the general online allocation problem. Building upon the effective greedy technique in the submodular optimization and the sequential allocating technique to guarantee fairness in Hosseini et al. (2024), we propose the Marginal-Greedy Algorithm (Algorithm 2), greedily assigning each arriving item to the agent who receives the maximum marginal value, for submodular binary valuation functions, and then show that the Marginal-Greedy Algorithm computes an NW allocation that is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW.
Input: An instance $(T,N,(v_{i})_{i\in N})$ with submodular binary valuation functions
Output: An NW approximate EF1 and approximate MMS allocation $\bf A$}
1 Initialize: let $\pi=(\pi_{1},\ldots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5 for $i=1$ to $n$ do
6
7    if $\Delta_{A_{\pi_{i}}}^{\pi_{i}}(e)>0$ then
8
9        Let $A_{\pi_{i}}=A_{\pi_{i}}\cup\{e\}$;
10       $\pi=(\pi_{1},\ldots,\pi_{i-1},\pi_{i+1},\ldots,\pi_{n},\pi_{i})$;
11        break;
12      end if
13
14   end for
15
16
17
return $\bf A$};
Algorithm 2 Marginal-Greedy Algorithm
In the analysis of results about submodular binary valuation functions, we effectively leverage submodularity and the properties of the matroid. There exists a relationship of submodularity and matroid: a function is submodular binary iff it is a rank function of a matroid Schrijver et al. (2003).
Definition 5.
A set system $\mathcal{M}=(E,\mathcal{F})$, where $E$ is a ground set and $\mathcal{F}$ is a family of subset of $E$, is a matroid if it satisfies the following properties: (1) $\emptyset\in\mathcal{F}$; (2) if $D\in\mathcal{F}$ and $C\subseteq D$, then $C\in\mathcal{F}$; (3) if $C,D\in\mathcal{F}$ and $|D|>|C|$, then there exists $x\in D\setminus C$ such that $C+x\in\mathcal{F}$. The rank function of matroid $\mathcal{M}$, denoted by $r(\cdot)$, is defined by: $r:2^{T}\rightarrow R$, $r(X)=\max\{|Y|:Y\subseteq T,Y\in\mathcal{F}\}$.
Theorem 2.
For the deterministic allocation of indivisible goods with submodular binary valuation functions, the Marginal-Greedy Algorithm (Algorithm 2) computes an NW allocation that satisfies $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS and $\frac{1}{2}$-max-USW.
Proof.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, where $v_{i}$ is submodular and binary for all $i\in N$. Fix an arbitrary round $k$, let $T^{k}=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ and $A^{k}_{0}$ the allocation and the set of all unallocated items at the end of round $k$ respectively.
In Algorithm 2, the arriving item is allocated to an agent who thinks that the marginal value of this item is $1$ if there exists such an agent. By Lemma 1, $\textbf{A}^{k}$ is NW. For each $v_{i}$, there exists a matroid $\mathcal{M}_{i}=(T^{k},\mathcal{F}_{i})$ such that $v_{i}$ is the rank function of $\mathcal{M}_{i}$. We have $v_{i}(A^{k}_{i})=|A^{k}_{i}|$, and $A^{k}_{i}\in\mathcal{F}_{i}$.
First, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-EF1. Consider any two agents $\in N, iji\neq j$. If $v_{i}(A^{k}_{j})>v_{i}(A^{k}_{i})$, by the property (3) of the matroid, there must exist an item $e\in A^{k}_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $B_{j}=\{e:e\in A^{k}_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$ and $B_{j}^{*}\in\arg\max\{|S|:S\subseteq B_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. We will show $|B_{j}|\leq|A^{k}_{i}|+1$. Assume that $|B_{j}|\geq|A^{k}_{i}|+2$, $j$ has the priority higher than $i$ at least $|B_{j}|\geq|A^{k}_{i}|+2$ times, which contradicts the order in the algorithm. Combining the property (3) of the matroid, we have $v_{i}(A^{k}_{j})=|A^{k}_{i}|+|B_{j}^{*}|\leq|A^{k}_{i}|+|B_{j}|\leq 2|A^{k}_{i}|+1$. Thus, ${\bf A}^{k}$ is $\frac{1}{2}$-EF1.
Second, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-MMS. Consider a MMS partition ${\bf X}=(X_{1},\ldots,X_{n})$ for agent $i$ of item set $T$, we assume that the share of agent $i$ in ${\bf A}^{k}$ is less than half of the maximin share of $i$, i.e. $v_{i}(A^{k}_{i})\leq\frac{1}{2}v_{i}(X_{j})$ for any $j\in N$. By the property (3) of the matroid, there exists an item $e\in X_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $C_{j}=\{e:e\in X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$ and $C_{j}^{*}\in\arg\max\{|S|:S\subseteq C_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. Then, we have $v_{i}(X_{j})=v_{i}(A^{k}_{i}\cup C_{j}^{*})=v_{i}(A^{k}_{i})+v_{i}(C_{j}^{*})=|A^{k}_{i}|+|C_{j}^{*}|$. Since $v_{i}(A^{k}_{i})=|A^{k}_{i}|\leq\frac{1}{2}v_{i}(X_{j})$ holds, we get $|C_{j}|\geq|C_{j}^{*}|>|A^{k}_{i}|+1$. Then, summing up all agents, we have $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. At last, we get $|\cup_{j\in N}C_{j}|=|\{e:e\in\cup_{j\in N}X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\{e:e\in T,A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\cup_{j\in N}B_{j}|\leq(n-1)(|A^{k}_{i}|+1)$, which contradicts the inequality $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. Thus, ${\bf A}^{k}$ is $\frac{1}{2}$-MMS.
Finally, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-USW. We make a slight adjustment to the Algorithm 2: when a new item arrives, if the marginal utility is 0 for all agents, then assign the item to agent $1$ without updating the agents’ order. Then, we will show that ${\bf A}^{k^{\prime}}=(A^{k}_{1}\cup A^{k}_{0},\ldots,A^{k}_{n})$ is the output of the adjusted algorithm. We only need to prove that in each round $s\in[k]$, if the marginal value of the arrived item $e_{s}$ to $A^{s-1}_{1}$ is 11, we have $v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1$. By the monotonicity of $v_{i}$, we get $v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0}\cup\{e_{s}\})\geq v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1=v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0})+1$, implying that the item $e_{s}$’s marginal value to $A^{s-1}_{1}\cup A^{s-1}_{0}$ is $1$. Thus, ${\bf A}^{k^{\prime}}=(A^{k}_{1}\cup A^{t}_{0},\ldots,A^{k}_{n})$ is the output of the adjusted algorithm. Since $v_{1}(A^{k}_{1}\cup A^{t}_{0})=v_{1}(A^{k}_{1})$, we have $USW({\bf A}^{k})=USW({\bf A}^{k^{\prime}})$. In other words, we only need to show that ${\bf A}^{k^{\prime}}$ is $\frac{1}{2}$-max-USW. The remaining proof part is based on the idea from Lehmann et al. (2001). Consider the first item $e_{1}$, we assume that $e_{1}$ is allocated to agent $i^{*}$ during the execution of the algorithm and $v_{i^{*}}(e_{1})=1$, $A_{i^{*}}^{1}=\{e_{1}\}$ and $A_{i}^{1}=\emptyset$ for $i\neq i^{*}$. Let ${\bf Y}=(Y_{1},\ldots,Y_{n})$ be an allocation achieving the optimal USW of instance $I=(T^{k},N,(v_{i})_{i\in N})$, and we assume that $e_{1}$ is allocated to agent $i^{o}$ in $Y$, $e_{1}\in Y_{i^{o}}$. Next, we construct a new instance ${I^{(1)}}=(T^{\prime},N,(v_{i}^{(1)})_{i\in N})$, where $v_{i^{*}}^{(1)}={\Delta^{i^{*}}_{e_{1}}}$ and $v_{i}^{(1)}=v_{i}$ for all $i\neq i^{*}$, $T^{\prime}=T^{k}\setminus\{e_{1}\}$, that is, the first item arriving is $e_{2}$, and then input the instance $I^{(1)}$ into the adjusted algorithm. Let initialized order of $N$ be $\pi=(\pi_{1},\ldots,\pi_{n},\pi_{i^{*}})$, the initialized allocation be ${\bf D}^{0}=(D^{0}_{1},\ldots,D^{0}_{n})$ , where $D_{i^{*}}=\{e_{1}\}$ and $D_{i}=\emptyset$ for all $i\neq i^{*}$. Assume that ${\bf D}^{k}=(D^{k}_{1},\ldots,D^{k}_{n})$ is the allocation at the end of round $k$. We have ${\bf D}^{k}={\bf A}^{k^{\prime}}$, then $USW({\bf D}^{k})=USW({\bf A}^{k^{\prime}})=USW({\bf A}^{k})$. Now, we construct an allocation ${\bf S}=(S_{1},\ldots,S_{n})$ for instance $I^{\prime}$ as follows: $S_{i^{o}}=Y_{i^{o}}\setminus\{e_{1}\}$, $S_{i}=Y_{i}$ for all $i\neq{i^{o}}$. It holds that $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})=v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$. By the monotonicity of $v_{i^{*}}$, we have $v_{i^{*}}^{(1)}(S_{i^{*}})=v_{i^{*}}(Y_{i^{*}}\cup\{e_{1}\})-v_{i^{*}}(\{e_{1}\})\geq v_{i^{*}}(Y_{i^{*}})-v_{i^{*}}(\{e_{1}\})$. $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$, where the first inequality follows from the submodularity and the second inequality follows from $v_{i^{o}}(\{e_{1}\})\leq v_{i^{*}}(\{e_{1}\})$. Let ${\bf Y}^{(1)}=(Y_{1}^{(1)},\ldots,Y_{n}^{(1)})$ be an allocation achieving the optimal USW of instance $I^{(1)}$. It holds that $USW({\bf Y}^{(1)})\geq USW({\bf S})=\cup_{i\in N}v_{i}^{(1)}(S_{i})=\cup_{i\in N,i\neq i^{*},i\neq i^{o}}v_{i}^{(1)}(S_{i})+v_{i^{o}}^{(1)}(S_{i^{o}})+v_{i^{*}}^{(1)}(S_{i^{*}})\geq USW({\bf Y})-2v_{i^{*}}(\{e_{1}\})$. By lemma 1 in Lehmann et al. (2001), $v_{i^{*}}^{(1)}$ is also submodular. So, by constructing the new instances $I^{(2)},\ldots,I^{(k-1)}$ and inducting each round based on the items that arrive, we have $USW({\bf A}^{k})=USW({\bf A}^{k^{\prime}})=USW({\bf D}^{k})=USW({\bf Y}^{k})\geq\frac{1}{2}USW({\bf Y}).$
Theorem 3.
For any $\epsilon>0$, with submodular binary valuation functions, no deterministic online algorithm can achieve
  • $(\frac{1}{2}+\epsilon)$-EF1 and non-wastefulness,
  • or $(\frac{1}{2}+\epsilon)$-MMS and non-wastefulness,
  • or $(\frac{1}{2}+\epsilon)$-max-USW and non-wastefulness.
Proof.
Figure 1: An instance with $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS allocations
First, let us consider EF1 or MMS. Consider an online allocation instance, where there are two agents $\{1,2\}$ with identical valuations $v$, and the online items belong to several categories $C_{1},\ldots,C_{l}$. In Figure 1, categories are distinguished by colours, and edges of different colours represent items of different categories. Denote a partition matroid $\mathcal{M}=(T,\mathcal{F})$, $\mathcal{F}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~k\in[l]\}$. The valuation $v$ is the rank function of the matroid $\mathcal{M}$; for any item set $S$, the value of $S$ is the number of items of the maximal independent set of $S$. For the first arrived item $e_{1}$, the marginal values for agent $1$ and agent $2$ are $1$. Without loss of generality, we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, in order not to break fairness, we must allocate it to agent $2$. For the third item $e_{3}$, the marginal values for agent $1$ and agent $2$ are $1$. We will discuss the following two cases.
Case 1: Allocate $e_{3}$ to agent $2$, shown in the left graph in Figure 1 Then, $A^{4}_{1}=\{e_{1}\}$$A^{4}_{2}=\{e_{2},e_{3},e_{4}\}$. We have that, $v(A^{4}_{1})=\frac{1}{2}v(A^{4}_{2}\setminus\{e\})$ for any $e\in A^{4}_{2}$, and $v(A^{4}_{1})=\frac{1}{2}\mathsf{MMS}_{1}$.
Case 2: Allocate $e_{3}$ to agent $1$, shown in the right graph in Figure 1; we must allocate the fourth item, $e_{4}$, to agent $1$ due to the requirement of non-wastefulness. Then, $A^{4}_{1}=\{e_{1},e_{3},e_{4}\}$$A^{4}_{2}=\{e_{2}\}$. We have that, $v(A^{4}_{2})=\frac{1}{2}v(A^{4}_{1}\setminus\{e\})$ for any $e\in A^{4}_{1}$, and $v(A^{4}_{2})=\frac{1}{2}\mathsf{MMS}_{2}$.
Next, let us consider USW. Consider an online allocation instance, where there are two agents $\{1,2\}$. For the first arrived item $e_{1}$, the marginal values for agents $1$ and $2$ are both $1$, $v_{1}(\{e_{1}\})=v_{2}(\{e_{1}\})=1$. Without loss of generality, we allocate $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, $v_{1}(\{e_{2}\})=v_{1}(\{e_{1},e_{2}\})=1$$v_{2}(\{e_{2}\})=0$, the marginal values for agents $1$ and $2$ are both 0. No matter which agent gets $e_{2}$, $USW(A^{2})=1$. However, the max-USW is $2$ by allocating $e_{1}$ to agent $2$ and $e_{1}$ to agent $1$. ∎
For an independent interest, we find that there is a side result of the Marginal-Greedy Algorithm, where it can also work for additive binary valuation functions to compute an NW allocation that is EF1, MMS, and max-USW, which demonstrates the robustness of our algorithm.
3.3 Additive Personalized Bi-valued Valuations
In this part, we consider the additive personalized bi-valued valuation function, which is often considered in the literature, such as Ebadian et al. (2022); Aziz et al. (2023). For any agent $i\in N$, her valuation function $v_{i}$ is additive personalized bi-valued if $v_{i}$ is additive and $v_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.
It is worth investigating whether an NW allocation satisfying EF1 or MMS is achievable with additive personalized bi-valued valuation functions, as the binary case is a special type of bi-valued function. However, we find that even for two agents with additive personalized bi-valued valuations, an NW allocation satisfying EF1 or MMS cannot be guaranteed.
3.3.1 Two Agents
We investigate the setting of additive personalized bi-valued valuations for two agents and establish the optimal non-trivial result for simultaneously approximating both EF1 and MMS. To achieve this objective, we propose the Adapted Envy-Graph Procedure (Algorithm 3), which requires sophisticated and subtle case analysis due to the uncertainty of future item values.
The Adapted Envy-Graph Procedure (AEGP) consists of three sub-algorithms: the Envy-Graph Procedure (EGP), the Preliminary-Breaking-Cycle (PBC), and the Deep-Breaking-Cycle (DBC), which are detailed in the Appendix. The core structure of our algorithm employs the EGP as the primary procedure. However, the EGP has a critical requirement: the input allocation is EF1 and must not contain any envy cycle.
When an envy cycle is detected, we execute the PBC and DBC algorithms as auxiliary procedures in subsequent rounds to allocate future items more effectively. Once the output allocation from either the PBC or DBC algorithm satisfies the EGP algorithm’s precondition, the EGP algorithm resumes execution. The execution flow of this integrated process is illustrated in Figure 2.
Figure 2: The execution flow of EGP, PBC, and DBC algorithms. The dashed line represents the execution of the corresponding algorithm. The solid line means that when some specific cases occur, we execute another algorithm.
The main challenge lies in addressing envy cycles effectively. This issue is straightforward to resolve in offline settings but becomes significantly more complex in online environments where reallocation is infeasible. To overcome this obstacle, we first introduce the PBC algorithm, which is developed through careful analysis of item values over the next two rounds. Specifically, the PBC algorithm can be executed consecutively for at most two rounds. After these two rounds, the PBC algorithm successfully addresses most cases, achieving EF1 allocation without an envy cycle. However, if the allocation remains be $\frac{1}{2}$-EF1, a more strategic approach is required in subsequent rounds to ensure fairer distribution. To address this limitation, we introduce the DBC algorithm. In the DBC algorithm, we establish allocation rules based on the parity of consecutive execution counts. Unlike the PBC algorithm, the DBC algorithm can be executed consecutively across multiple rounds while maintaining the $\frac{1}{2}$-EF1 approximation guarantee, before its output allocation satisfies the precondition of the EGP. In summary, both the PBC and DBC algorithms establish specific allocation rules for future items when an envy cycle emerges. These rules not only ensure EF1 approximation throughout their execution but also continuously monitor whether the EGP precondition can be satisfied and actively work to achieve it when feasible.
Definition 6 (Envy Graph for Goods).
Given an allocation ${\bf A}^{k}$, the corresponding envy graph is defined as $G=(V,E)$, where the vertex set $V$ corresponds to agent set $N$, and a directed edge $(i,j)\in E$ iff agent $i$ envies agent $j$, i.e., $v_{i}(A_{i}^{k})<v_{i}(A_{j}^{k})$. Additionally, the directed cycle in the envy graph is called an envy cycle.
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive personalized bi-valued valuation functions
Output: An NW approximate EF1 and approximate MMS allocation $\bf A$}
1
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
Initialize $\alpha=1$; // Algorithm Indicator
3
Initialize $(\lambda_{1},\mu_{1})$ and $(\lambda_{2},\mu_{2})$, where $\lambda_{1}=\lambda_{2}=1$ and $\mu_{1}=\mu_{2}=0$; ;
  // $\lambda$: the times of consecutive execution for the PBC or DBC algorithm, and $\mu$: the case label in the PBC or DBC algorithm.
4
5Build an envy graph $G$ for two agents;
6 when item $e\in T$ arrives do
7 if $\alpha=1$ then
8
9    ${\bf A},\alpha\leftarrow\mathsf{EGP}(N,{\bf A},\alpha,e)$;
10    if $\alpha=2$ then
       ${\bf A},\alpha\leftarrow\mathsf{PBC}(N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1}))$ ;
          // Envy cycle will appear.
11
12      end if
13
14 else if $\alpha=2$ then
15    ${\bf A},\alpha,(\lambda_{1},\mu_{1})\leftarrow\mathsf{PBC}(N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1}))$;
16 else if $\alpha=3$ then
17    ${\bf A},\alpha,(\lambda_{2},\mu_{2})\leftarrow\mathsf{DBC}(N,{\bf A},\alpha,e,(\lambda_{2},\mu_{2}))$;
18   end if
19
20  Update the envy graph $G$;
21
return $\bf A$};
Algorithm 3 Adapted Envy-Graph Procedure
Theorem 4.
For the deterministic allocation of indivisible goods for two agents with additive personalized bi-valued valuations, the AEGP algorithm (Algorithm 3) computes an NW allocation that satisfies $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
The following two lemmas are crucial in the proof of Theorem 4.
Input: $N,A,\alpha,e$
Output: An NW allocation A and $\alpha$
1
2if there is no envy between the two agents  then
3 if $\exists i\in N$, $v_{i}(e)=b_{i}$  then
4      Choosing the agent $i$ who values $e$ at $b_{i}$ (breaking ties arbitrarily);
5 else
6      Arbitrarily choose one agent $i$;
7   end if
8
9else
10   Let $i\in N$ be an unenvied agent and j=2ij=2-i;
11 if $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ then
12    $\alpha=2$;
13    return A and $\alpha$;
14   end if
15
16 end if
[…]
return A and $\alpha$;
Algorithm 4 EGP (Envy-Graph Procedure)
Lemma 3.
The EGP algorithm (Algorithm 4) computes an NW allocation that is EF1 without an envy cycle.
Proof.
Note that there are three kinds of input allocation. The first one is the initial empty allocation, which is EF1 without an envy cycle. The second one is the output allocation of the PBC algorithm. The last one is the output allocation of the DBC algorithm. It suffices to show that the last two kinds of input allocation are EF1 without an envy cycle. That is because if the input allocation of the EGP algorithm is EF1 without an envy cycle, by the description of the algorithm and induction, it is easy to check that the property of EF1 without an envy cycle is always maintained. By Lemma 4, we can derive that when the PBC or DBC algorithm outputs $\alpha=1$, implying that the EGP algorithm is ready to be executed, the output allocation of the PBC or DBC algorithm is EF1 without an envy cycle. Therefore, our claim holds. ∎
Lemma 4.
The PBC or DBC algorithm (Algorithm 5 or 6) computes an NW allocation satisfying
  • $\frac{1}{2}$-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
We split the proof of Lemma 4 into two parts, where we show that the claim holds for both the PBC and DBC algorithms.
Observation 1.
When the PBC algorithm (Algorithm 5) is executed, for the first arriving item $e$, we have $v_{i}(e)=a_{i}$.
Proof.
The reason why the PBC algorithm is executed is that for the arriving item $e$, if we add it to agent $i$’s bundle, there will be an envy cycle between two agents. i.e., $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ and $v_{j}(A_{j})<v_{i}(A_{i}\cup\{e\})$. Suppose for the contradiction that $v_{i}(e)=b_{i}$. Since the input allocation $A$ is EF1, for agent $i$, we have $v_{i}(A_{j})-v_{i}(A_{i})\leq b_{i}$. Thus, allocating item $e$ to agent $i$ eliminates the envy and cannot lead to an envy cycle, which is a contradiction. Therefore, for the arriving item $e$, it holds that $v_{i}(e)=a_{i}$. ∎
Observation 2.
When the PBC algorithm (Algorithm 5) is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
Proof.
Suppose that $|A_{i}|=0$. In that case, we have $|A_{j}|=1$. Otherwise, there would be an envy cycle. After allocating $e$, there will be an envy cycle, i.e., $v_{i}(e)=a_{i}<v_{i}(A_{j})=b_{i}$ and $v_{j}(A_{j})<v_{j}(e)$. If the second inequality holds, we have $v_{j}(A_{j})=a_{j}$ but $v_{i}(A_{j})=b_{i}$, which contradicts the line 4 in the EGP algorithm (Algorithm 4). Suppose that $|A_{j}|=0$. In that case, we have $|A_{i}|=0$, which contradicts the fact that after allocating $e$, an envy cycle appears. ∎
Lemma 5.
The PBC algorithm (Algorithm 5) computes an NW allocation satisfying:
  • $\frac{1}{2}$-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
Proof.
When the PBC algorithm is implemented ($\alpha=2$), the input allocation $\textbf{A}=(A_{1},A_{2})$ output by the EGP algorithm (Algorithm 4) is EF1 without an envy cycle. By the description of the PBC algorithm, it is executed in at most two consecutive rounds.
$\lambda_{1}=1$. For the first arriving item $e_{1}$, by Observation 1, we have $v_{i}(e_{1})=a_{i}$. If $\lvert A_{j}\rvert=1$, we have $v_{i}(A_{j})=b_{i}$ since agent $i$’s bundle is EF1. Then, agent $j$ picks this item. For agent $i$, her bundle is EF1 since we have $\lvert A_{i}\lvert\geq 1$ and $v_{i}(e_{1})=a_{i}$. For agent $j$, her new bundle is EF since she does not envy agent $i$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $\lvert A_{j}\rvert>1$, let us consider two cases about the valuation of item $e_{1}$.
  • If $v_{j}(e_{1})=b_{j}$, item $e$ is allocated to agent $j$. Then, we set $\lambda_{1}=2$ and $\mu_{1}=1$. In this case, for agent $j$, her new bundle is EF since $v_{j}(A_{j}\cup\{e_{1}\})>v_{j}(A_{j})\geq v_{j}(A_{i})$ and for agent $i$, her bundle is $\frac{1}{2}$-EF1, i.e, $v_{i}(A_{i})\geq\frac{1}{2}v_{i}(A_{j}\cup\{e_{1}\}\setminus\{g\})$ holds for $g\in A_{j}\cup\{e_{1}\}$, since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(A_{i})\geq a_{i}=v_{i}(e_{1})$. Thus, this allocation is $\frac{1}{2}$-EF1.
  • If $v_{j}(e_{1})=a_{j}$, item $e$ is allocated to agent $i$. Then, we set $\lambda_{1}=2$ and $\mu_{1}=2$. In this case, for agent $i$, her new bundle is still EF1 since she is an unenvied agent in Algorithm 3 and the input allocation $A$ is EF1, and for agent $j$, her bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ but $v_{j}(A_{j})<v_{j}(A_{j}\cup\{e_{1}\})$. Thus, this allocation is EF1 with an envy cycle.
It is clear that any possible allocation from the above cases does not satisfy EF1 without an envy cycle. Thus, we keep $\alpha=2$.
$\lambda_{1}=2$. The second item $e_{2}$ arrives. Let us consider two cases about $\mu_{1}$.
  • If $\mu_{1}=1$, agent $i$ picks it. In this case, for agent $j$, her bundle is still EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. For agent $i$, her new bundle is EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{2})\geq v_{i}(e_{1})=a_{i}$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$.
  • If $\mu_{1}=2$, let us consider two cases. When $v_{i}(e_{2})=a_{i}$, agent $j$ picks it. For agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{1})=v_{i}(e_{2})=a_{i}$. For agent $j$, her new bundle is EF because $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{2})=v_{j}(e_{1})=a_{j}$. Hence, the allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$. When $v_{i}(e_{2})=b_{i}$, it is allocated to agent $i$. For agent $i$, her new bundle is EF since it holds that $v_{i}(A_{i}\cup\{e_{1},e_{2}\})=v_{i}(A_{i})+a_{i}+b_{i}\geq v_{i}(A_{j}\setminus\{g\})+v_{i}(g)+a_{i}$ for some $g\in A_{j}$ where the inequality follows from the fact that $v_{i}(g)\in\{a_{i},b_{i}\}$. For agent $j$, her bundle is $\frac{1}{2}$-EF1, i.e., $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\cup\{e_{1},e_{2}\}\setminus\{g\})$ holds for some $e\in A_{i}\cup\{e_{1},e_{2}\}$, since $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(A_{j})\geq v_{j}(e_{1})=\min\{v_{j}(e_{1}),v_{j}(e_{2})\}$. Thus, this allocation is $\frac{1}{2}$-EF1, and we set $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$.
Combining the above cases, it is not hard to see that our claim is correct. ∎
Observation 3.
When the DBC algorithm (Algorithm 6) is executed, there are two properties about input allocation $\bf A$:
  • $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$;
  • Let $e_{1}$ and $e_{2}$ denote two new allocated items in the PBC algorithm, we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$, where $e\in\arg\max_{g\in\{e_{1},e_{2}\}}v_{j}(g)$.
Proof.
When the DBC algorithm is ready to be executed, the PBC algorithm is terminated for $\lambda_{1}=\mu_{1}=2$. Then, by the proof of Lemma 5, it is easy to see that the statements are correct. ∎
Lemma 6.
The DBC algorithm (Algorithm 6) computes an NW allocation satisfying
  • $\frac{1}{2}$-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
Proof.
By Observation 3, the input allocation $\textbf{A}=(A_{1},A_{2})$ is $\frac{1}{2}$-EF1, where $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$ for some $e\in A_{i}$. Unlike the PBC algorithm, this algorithm may be executed in many rounds. Let $g_{1}$ and $g_{2}$ denote the two online items that are allocated to agent $i$ in the PBC algorithm and we have $\min\{v_{j}(g_{1}),v_{j}(g_{2})\}=a_{j}$. Currently, we have $\alpha=3$.
$\lambda_{2}=1$. For the arriving item $e_{1}$ when $\lambda_{2}=1$, let us consider three cases of agents’ valuations.
  • When $v_{i}(e_{1})=a_{i}$, agent $j$ picks it. There are two subcases about $v_{j}(e_{1})$. If $v_{j}(e_{1})=b_{j}$, for agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an envy cycle. If $v_{j}(e_{1})=a_{j}$, for agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=a_{i}=\min\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an envy cycle. Combining the above two cases, we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=1$, implying that the EGP algorithm is ready to be executed.
  • When $v_{i}(e_{1})=b_{i}$ and $v_{j}(e_{1})=b_{j}$, it is allocated to agent $j$ and we set $\mu_{2}=1$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is also EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 with an envy cycle, and we set $\lambda_{2}=2$.
[…]
$\lambda_{2}=2$. For the arriving item $e_{2}$ when $\lambda_{2}=2$, let us consider two cases about $\mu_{2}$.
  • When $\mu_{2}=1$, we need to consider two subcases. If $v_{i}(e_{2})=b_{i}$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{2})=b_{i}=v_{i}(e_{1})$. For agent $j$, her bundle is $\frac{1}{2}$-EF1 since we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda_{2}=1$, $\mu_{2}=0$, and keep $\alpha=3$, implying that the DBC algorithm continues to be executed. If $v_{i}(e_{2})=a_{i}$, it is allocated to agent $j$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$,we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{2}\})$. For agent $j$, her new bundle is trivially EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq a_{i}=v_{j}(g_{1})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$, implying the EGP algorithm is ready to be executed.
  • When $\mu=2$, it is directly allocated to agent $j$. For agent $i$, her bundle is EF since $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{1})=b_{i}\geq v_{i}(e_{2})$. For agent $j$, her new bundle is $\frac{1}{2}$-EF1 since $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{i}$ and $v_{j}(e_{2})\geq v_{j}(e_{1})=a_{j}$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda_{2}=1$, $\mu_{2}=0$, and keep $\alpha=3$, implying that the DBC algorithm continues to be executed.
Combining the above cases, our claim holds. ∎
Input: $N,A,\alpha,e,(\lambda_{1},\mu_{1})$
Output: An NW allocation A, $\alpha$, and $(\lambda_{1},\mu_{1})$
1
2if $\lambda_{1}=1$ then
3
4 if $\lvert A_{j}\rvert=1$ then
5
6    $A_{j}=A_{j}\cup\{e\}$;
7    $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
8 else
9
10    if $v_{j}(e)=b_{j}$ then
11
12       $A_{j}=A_{j}\cup\{e\}$;
13       $\lambda_{1}=2$ and $\mu_{1}=1$;
14    else
15       $A_{i}=A_{i}\cup\{e\}$;
16       $\lambda_{1}=2$ and $\mu_{1}=2$;
17      end if
18
19   end if
20
21else
22
23 if $\mu_{1}=1$  then
24
25    $A_{i}=A_{i}\cup\{e\}$;
26 else
27
28    if $v_{i}(e)=a_{i}$ then
29
30       $A_{j}=A_{j}\cup\{e\}$;
31    else
32
33       $A_{i}=A_{i}\cup\{e\}$;
34       $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$;
35      end if
36
37   end if
38
39 $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
40 end if
41
return A, $\alpha$ and $(\lambda_{1},\mu_{1})$;
Algorithm 5 PBC (Preliminary-Breaking-Cycle)
Input: $N,A,\alpha,e,(\lambda_{2},\mu_{2})$
Output: An NW allocation A, $\alpha$, and $(\lambda_{2},\mu_{2})$
1
2if $\lambda_{2}\mod 2=1$ then
3
4 if $v_{i}(e)=a_{i}$ then
5    $A_{j}=A_{j}\cup\{e\}$;
6    $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
7 else if $v_{i}(e)=b_{i}$ and $v_{j}(e)=b_{j}$ then
8
9    $A_{j}=A_{j}\cup\{e\}$;
10    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=1$;
11 else
12    $A_{i}=A_{i}\cup\{e\}$;
13    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=2$;
14   end if
15
16 else if $\lambda_{2}\mod 2=0$ then
17 if  $\mu_{2}=1$  then
18    if $v_{i}(e)=b_{i}$ then
19
20       $A_{i}=A_{i}\cup\{e\}$;
21       $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
22    else
23       $A_{j}=A_{j}\cup\{e\}$;
24       $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
25      end if
26
27 else
28
29    $A_{j}=A_{j}\cup\{e\}$;
30    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
31   end if
32
33 end if
34return A, $\alpha$, $(\lambda_{2},\mu_{2})$;
Algorithm 6 DBC (Deep-Breaking-Cycle)
Proof of Theorem 4. In the AEGP, it can be seen that for different values of the algorithm indicator $\alpha$, we need to choose the corresponding algorithm, i.e., $\alpha=1$: EGP, $\alpha=2$: PBC, or $\alpha=3$: DBC, to execute. By Lemmas 3 and 4, it can be seen that the output allocation is NW and $\frac{1}{2}$-EF1. Regarding MMS, by the implication between EF1 and MMS Amanatidis et al. (2018), i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{\alpha}{(n-1)\alpha+1}$-MMS, where $n$ is the number of agents, it is clear that the output allocation is $\frac{\frac{1}{2}}{\frac{1}{2}+1}=\frac{1}{3}$-MMS.∎
For an independent interest, in the Appendix, we show that if each online item has a deadline of one period, i.e., each arriving item can wait for at most one round to be allocated, EF1 allocation can be found in the above setting. At last, we present the impossible results to show that our approximation guarantees for EF1 and MMS are tight, and the approximate EF1 (or MMS) is incompatible with USW.
Theorem 5.
For two agents with additive personalized bi-valued valuation functions, no deterministic online algorithm can compute an NW allocation that is $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{3}+\epsilon)$-MMS for any $\epsilon>0$.
Proof.
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 $\epsilon$ 11 $\epsilon$ 1
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1
Case 1
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
agent 11 $\epsilon$ 11 $\epsilon$ $\epsilon$ 11
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1 11
Case 2-1
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 $\epsilon$ 11 $\epsilon$ $\epsilon$
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1
Case 2-2
Table 6: The impossibility results of EF1 and MMS for additive personalized bi-valued valuations for two agents.
Let $a_{1}=a_{2}=\epsilon$ and $b_{1}=b_{2}=1$, where $0<\epsilon<1$. Consider the following online instance for $n=2$, where the value of each item is shown in Table 6. When $t=1$, the item with the value of $\epsilon$ for both agents arrives. Without loss of generality, agent 1 picks it. When $t=2$, the arriving item has the value of 1 for agent 1 and $\epsilon$ for agent 2. If it is allocated to agent 1, agent 2’s bundle is empty, resulting in 0-EF1 and 0-MMS. Thus, it should be allocated to agent 2. Then, when $t=3$, the arriving item has the value of $\epsilon$ for both agents. Let us consider two different choices of this item.
Case 1. The arriving item in round $t=3$ is allocated to agent 1. Then, in round $t=4$, the item that has the value of 1 for both agents arrives. Since we consider non-wastefulness, one agent must pick it. If it is allocated to agent 1, for agent 1, her bundle is EF1 and MMS to her, and for agent 2, her bundle is $\frac{\epsilon}{1+\epsilon}$-EF1 and $\frac{1}{3}$-MMS. If it is allocated to agent 2, for agent 1, her bundle is $2\epsilon$-EF1 and $\frac{2\epsilon}{1+\epsilon}$-MMS, and for agent 2, her bundle is EF1 and MMS. Therefore, we can derive that the upper bound of EF1 is $2\epsilon$ and MMS is $\frac{1}{3}$ respectively.
Case 2. The arriving item in round $t=3$ is allocated to agent 2. Then, in round $t=4$, the item has the value of $\epsilon$ for agent 1 and 1 for agent 2. Let us consider two choices for allocating this item.
  • If it is allocated to agent 1, this allocation is EF1 and $\frac{2}{3}$-MMS. At last, in round $t=5$, the arriving item has the value of 1 for both agents. To guarantee non-wastefulness, it must be allocated to some agent. It is not hard to check that no matter which agent picks it, the final allocation is $\frac{2\epsilon}{1+\epsilon}$-EF1 and $\frac{2\epsilon}{1+\epsilon}$-MMS.
  • If it is allocated to agent 2, there is no further arriving item. For agent 1, her bundle is $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS, and for agent 2, her bundle is EF and MMS. Thus, this allocation is $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
Therefore, we can derive that the upper bound of EF1 is $\frac{1}{2}$ and MMS is $\frac{1}{3}$ respectively. Combining the above two cases, it is clear that no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\beta$-MMS for any $\alpha>\frac{1}{2}$ and $\beta>\frac{1}{3}$. ∎
Proposition 2.
Given any $\alpha,\beta>0$, for the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation satisfying
  • $\alpha$-EF1 and $\beta$-max-USW,
  • or $\alpha$-MMS and $\beta$-max-USW,
even for two agents with additive personalized bi-valued valuations.
Proof.
Without loss of generality, we assume that $v_{i}\in\{1,\frac{1}{\epsilon}\}$ and $v_{2}\in\{\epsilon^{2},\epsilon\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{1+\frac{1}{\epsilon}}=\frac{\epsilon^{3}+\epsilon}{1+\epsilon}$-max-USW. For any $\beta>0$, let $\epsilon<\beta^{2}$, we can derive that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 (or $\alpha$-MMS) and $\beta$-max-USW for any $\alpha,\beta>0$
3.3.2 The Combination with Additive Binary Valuations
In this part, we focus on the setting where only one agent (agent $n$) has an additive personalized bi-valued valuation function, while all other agents share identical additive binary valuation functions, formed by $(T,N,(v_{i})_{i\in N\setminus\{n\}},v_{n})$. We show that an NW allocation that is EF1 and MMS always exists and can be computed by the Adapted-Picking Algorithm (Algorithm 7). The high-level idea of our algorithm is to allocate the arriving items that all agents value at non-zero values “evenly”, according to the requirement that agent $n$ has the top priority of picking the item with the value of $b_{n}$ and the lowest priority of picking the item with the value of $a_{n}$. In that case, agent $n$ takes advantage of the item options, and then we can guarantee that EF1 and MMS hold simultaneously. In Table 7, we illustrate how our algorithm works for two agents.
Input: An instance $(T,N,(v_{i})_{i\in N\setminus\{n\}},v_{n})$ with identical (n1)(n-1) additive binary valuation functions $(v_{1},\ldots,v_{n})$ and one additive personalized bi-valued valuation function $v_{n}$
Output: An NW, EF1, and MMS allocation $\bf A$}
1
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5  Let $i_{min}\in\arg\min_{i\in[n-1]}|A_{i}|$;
6 if $v_{i_{min}}(e)=0$ then
7
8    $A_{n}=A_{n}\cup\{e\}$;
9 else
10
11    if $v_{n}(e)=b_{n}$ then
12
13       if $v_{i_{min}}(A_{n})\leq v_{i_{min}}(A_{i_{min}})$ then
14          $A_{n}=A_{n}\cup\{e\}$;
15       else
16
17          $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
18         end if
19
20    else
21       if $v_{i_{min}}(A_{n})>v_{i_{min}}(A_{i_{min}})$ then
22
23          $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
24       else
25          $A_{n}=A_{n}\cup\{e\}$;
26         end if
27
28      end if
29
30   end if
31
32
return $\bf A$};
Algorithm 7 Adapted-Picking Algorithm
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $\ldots$
agent 11 11 11 0 1 1 $\ldots$
agent 2 $a_{2}$ $a_{2}$ $a_{2}$ $b_{2}$ $a_{2}$ $\ldots$
Table 7: A simple illustration for two agents of Algorithm 7.
Theorem 6.
For the deterministic allocation of indivisible goods for (n1)(n-1) agents with identical additive binary valuation functions and one agent with the additive personalized bi-valued valuation function, the Adapted-Picking Algorithm (Algorithm 7) computes an NW, EF1, and MMS allocation.
Proof.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$, the allocation at the end of round $t$ in Algorithm 7.
EF1.
We show that the allocation in any round is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after allocating item $e_{k}$, the allocation $\textbf{A}^{k}$ is EF1. Next, we show that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k}$ is still EF1. When item $e_{k+1}$ arrives, let us consider several cases about the allocation of this item.
  • If $v_{i_{min}}(e_{k+1})=0$, it is allocated to agent $n$. For any agent $i\in[n-1]$, she values it at zero, so her bundle is still EF1 to her. For agent $n$, the value of her bundle increases, so her bundle is also EF1 to her.
[…]
Combining the above cases, it can be concluded that in round $k+1$, the allocation is EF1. This completes the induction and establishes the correctness of our claim.
MMS.
Fix an arbitrary round $k$. Pick an arbitrary agent $i\in[n-1]$. For agent $i$, assume that the number of items with the value of one that arrive between round $t=1$ and $k$ is $s_{i}^{k}$. By the description of our algorithm, items are almost evenly allocated to agents with identical additive binary valuations. Thus, it is trivial that $v_{i}(A_{i})=|v_{i}(A_{i})|\geq\mathsf{MMS}_{i}=\lfloor\frac{s_{i}^{k}}{n}\rfloor$. For agent $n$, if she picks item $e_{k}$, by the above proof of EF1, her bundle is EF, implying that her bundle is also MMS; If she does not pick item $e_{k}$, by the above proof of EF1, the value of her bundle is at most $a_{n}$ less than that of the envied agent’s bundle. Considering that she has the priority of picking the item with the value of $b_{n}$ and the number of items in her bundle is at most one less than that of the other agent’s bundle, it is not hard to check that her bundle is still MMS. ∎
At last, we complement the above result by showing the incompatibility of EF1 (or MMS) and USW.
Proposition 3.
Given any $\alpha>0$, for the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation satisfying
  • EF1 and $\alpha$-max-USW,
  • or MMS and $\alpha$-max-USW,
even for two agents, where one agent has the additive binary valuation function, and the other agent has the additive personalized bi-valued valuation function.
Proof.
Without loss of generality, we assume that $v_{1}\in\{0,1\}$ and $v_{2}\in\{\epsilon^{2},\epsilon\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{2}\approx\frac{1}{2}$-max-USW. When $t=3$, the item $e_{3}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 1, this allocation is $\epsilon$-EF1, $\epsilon$-MMS, and $\frac{2+\epsilon^{2}}{3}\approx\frac{2}{3}$-max-USW. If it is allocated to agent 2, this allocation is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1+\epsilon^{2}+\epsilon}{3}\approx\frac{1}{3}$-max-USW. In other words, no allocation is EF1 or MMS. Combining the above cases, it can be seen that if an online algorithm outputs an NW allocation that is EF1 or MMS in the setting of additive binary and additive personalized bi-valued valuations, it may not guarantee $\alpha$-max-USW for any given $\alpha>0$. ∎
4 Allocation of Chores
In this section, we study the setting of indivisible chores. Because “online” chores can be regarded as daily compulsory tasks or duties, it is unreasonable to throw them away. In other words, for each arriving chore, there must be some agent taking it. Therefore, we first introduce the following constraint.
Definition 7 (Completeness).
An allocation ${\bf A}^{k}$ is complete if, for any chore $e$ that has arrived by round $k$, it is not discarded.
Here are the fairness and efficiency definitions used in the chores setting.
Definition 8 (Envy-freeness up to One Chore).
For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate envy-free up to one chore ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i}^{k}\neq\emptyset$, it holds that $c_{i}(A_{i}^{k}\setminus\{e\})\leq\alpha\cdot c_{i}(A_{i^{\prime}}^{k})$ for some $e\in A_{i}^{k}$.
Definition 9 (Minimax Share Fairness).
For any agent $i$, her minimax share $\mathsf{MMS}_{i}^{k}$ by round $k$ is defined as:
$\mathsf{MMS}_{i}^{k}=\min_{{\bf B}^{k}\in\Pi_{n}(T^{k})}\max_{i^{\prime}\in N}c_{i}(B_{i^{\prime}}^{k}),$
where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$. For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate minimiax share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $c_{i}(A_{i}^{k})\leq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
Definition 10 (Utilitarian Social Cost).
The utilitarian social cost $USC$ of allocation ${\bf A}^{k}$ is given by $USC({\bf A}^{k})=\sum_{i\in N}c_{i}(A_{i}^{k})$. For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-min-USC if it holds that $USC({\bf A}^{k})\leq\alpha\cdot\min_{{\bf B}^{k}\in\Pi_{n}(T^{k})}USC({\bf B}^{k})$, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
4.1 The Impossibility of Approximate EF1
In this part, we show the following negative result about the approximation of EF1 for additive personalized tri-valued cost functions. For any agent $i\in N$, her cost function $c_{i}$ is additive personalized tri-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
Theorem 7.
No deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$, even for two agents with additive personalized tri-valued cost functions.
Proof.
$e_{1}$ $e_{2}$ $e_{3}$
agent 11 11 $\epsilon$ $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Table 8: The impossibility result of EF1 for additive personalized tri-valued cost functions.
Let $a_{1}=a_{2}=\epsilon$, $b_{1}=b_{2}=1$, and $z_{1}=z_{2}=\frac{1}{\epsilon}$, where $0<\epsilon<1$. Consider the following online instance for $n=2$, where the cost of each item is shown in Table 8. When $t=1$, the arriving item has the cost of 1 for both agents. Without loss of generality, agent 1 picks it. When $t=2$, the item that has the cost of $\epsilon$ for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. In order to avoid the approximation ratio of EF1 being infinite, this item must be allocated to agent 2. At last, in round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. In this case, if this item is allocated to agent 1, for agent 2, the approximation ratio of EF1 is $\frac{1}{\epsilon}$; if this item is allocated to agent 2, for agent 1, the approximation ratio of EF1 is $\frac{1}{\epsilon}$. Therefore, for any $\alpha>0$, let $\epsilon<\frac{1}{\alpha}$, and we can conclude that no algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any given $\alpha>0$. This impossibility result can be generalized to more than two agents, and we defer the proof to Proposition 4. ∎
Proposition 4.
For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$ for more than two agents with additive personalized tri-valued cost functions.
Proof.
$e_{1}$ $e_{2}$ $\ldots$ $e_{n-2}$ $e_{n-1}$ $e_{n}$ $e_{n+1}$
agent 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent 22 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent n2n-2 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent n1n-1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Table 9: The impossibility result of EF1 for additive personalized tri-valued cost functions when there are more than two agents.
We extend the impossibility result of approximate EF1 for chores to $n\geq 3$ agents. Consider the following instance with $n\geq 3$ agents, where the cost of each item is shown in Table 9. Before each agent’s bundle has one agent, we should allocate the arriving item to the agent with an empty bundle to avoid the approximation ratio of EF1 being infinite. For $t\leq n-2$, each arriving item has the cost of $\frac{1}{\epsilon}$ for each agent. Without loss of generality, assume that agent $i\in[n-2]$ picks the item that arrives at round $i$. When t=n1t=n-1, the arriving item has the cost of 1 for all agents. Assume that it is allocated to agent n1n-1. When $t=n$, the arriving item has the cost of $\frac{1}{\epsilon}$ for agent $n$ and $\epsilon$ for the remaining agents. It is clear that agent $n$ picks it. In the last round $t=n+1$, the item that has the cost of $\frac{1}{\epsilon}$ for all agents arrives. If it is allocated to agent $i\in[n-2]$, we can derive that the approximation ratio of EF1 is $\frac{1}{\epsilon}$. If it is allocated to agent n1n-1 or $n$, the allocation is also $\frac{1}{\epsilon}$-EF1. Thus, the allocation in this round is $\frac{1}{\epsilon}$-EF1. Given any $\alpha>1$, let $\epsilon>\frac{1}{\alpha}$, we can conclude that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha>1$. ∎
By the above theorem, no approximation of EF1 can be guaranteed for additive personalized tri-valued cost functions. Like the negative result of goods setting, we cannot make further progress in the approximation of EF1 beyond additive personalized tri-valued cost functions if there is no prior information about future items. Therefore, in the following parts, we focus on additive binary, supermodular binary, and additive personalized bi-valued cost functions, and study whether there are positive results about the approximation of EF1.
4.2 Additive/Supermodular Binary Costs
In this part, we consider (additive) binary and supermodular binary cost functions in the chores setting.
4.2.1 Additive Binary Costs
For additive binary cost functions, where for any agent $i\in N$, $c_{i}$ is additive and $c_{i}(e)\in\{0,1\}$ for any $e\in T$, we propose the Compelled-Greedy Algorithm (Algorithm 8) to compute a complete allocation that satisfies EF1, MMS, and min-USC.
Input: An instance $(T,N,(c)_{i\in N})$ with additive binary cost functions
Output: An complete EF1 allocation $\bf A$}
1 Initialize: let $\pi=(\pi_{1},\ldots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5 for $i=1$ to $n$ do
6
7    if $v_{\pi_{i}}(\{e\})=0$ then
8         Let $A_{\pi_{i}}=A_{\pi_{i}}\cup\{e\}$;
9        break;
10      end if
11
12   end for
13 $A_{\pi_{1}}=A_{\pi_{1}}\cup\{e\}$;
14 $\pi=(\pi_{2},\ldots,\pi_{i-1},\pi_{i+1},\ldots,\pi_{n},\pi_{1})$;
15
16
17return $\bf A$};
Algorithm 8 Compelled-Greedy Algorithm
Theorem 8.
For additive binary cost functions, the Compelled-Greedy Algorithm (Algorithm 8) computes a complete allocation that satisfies EF1, MMS, and min-USC.
Proof.
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, where $c_{i}$ is additive and binary for all $i\in N$. Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A^{k}}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $k$. For every item $e\in T$, if there exists one agent such that the cost about $e$ of some agent is 0, the algorithm must allocate $e$ to some agent $a$ with $v_{a}(e)=0$. Otherwise, the costs of all agents are $1$. Regardless of how it is allocated, the increment brought to utilitarian social welfare is $1$, so $A^{k}$ is a min-USC allocation.
Next, we show that $\bf A^{k}$} is EF1. Assume that there exists $i,j\in N$ such that, for any item $e\in A^{k}_{i}$,
$c_{i}(A^{k}_{j})\leq c_{i}(A^{k}_{i}\setminus\{e\}).$ (1)
There must exist an item $e^{i}\in A^{k}_{i}$ such that $c_{i}(A^{k}_{i})-c_{i}(A^{k}_{i}\setminus\{e\})=1$. Without loss of generality, we assume $e^{i}$ is the last item with $1$ cost, and it arrives at $\ell$-th round. For any agent, the cost of $e^{i}$ is $1$, and agent $i$’s priority is higher than $j$. Due to the algorithm’s order in each round, the number of items with a $1$ cost of agent $j$ must be at least as large as the number of items with a $1$ cost of agent $j$. It contradicts inequality (1).
Finally, we show that $\bf A^{k}$} is MMS. For any agent $j\in N$ and any item $e\in A^{k}_{j}$, if the cost of $c_{j}(\{e\})$ is $1$, $c_{i}(\{e\})=1$ for all $i\in N$. So, $c_{i}(A^{r}_{j})\geq c_{j}(A^{r}_{j})$ for any $r$-th round. $e^{i}$ is the last item with $1$ cost and it arrives at $\ell$-th round, let $T_{\ell}$ be the set arrived after $\ell$-th round. In the $\ell-1$-th round, $c_{i}(A^{\ell}_{i})=c_{i}(A^{\ell-1}_{i})+1\leq\frac{\sum_{j\in N}c_{i}(A^{\ell-1}_{j})}{n}+1=\frac{c_{i}(T_{\ell-1})}{n}+1$. Since $c_{i}(A^{\ell}_{i})$ is an integer, we have $c_{i}(A^{k}_{i})=c_{i}(A^{\ell}_{i})\leq\lceil\frac{c_{i}(T_{\ell})}{n}\rceil\leq\lceil\frac{c_{i}(T)}{n}\rceil=\mathsf{MMS}_{i}$. ∎
4.2.2 Supermodular Binary Costs
For supermodular binary cost functions, where for any agent $i\in N$, $c_{i}$ is supermodular and the marginal cost of any item is 0 or 1, we first obtain a significant negative result regarding the approximation of fairness and efficiency. This contrasts starkly with the findings in offline settings, where it is possible to efficiently obtain an allocation satisfying EF1, MMS, or min-USC Barman et al. (2023).
Lemma 7.
(Lemma 2 in Barman et al. (2023)) A cost function $c$ is supermodular and binary if and only if the function $g$, g(S)=|S|c(S)g(S)=|S|-c(S) for any set $S$, is a matroid rank function.
Theorem 9.
For any $\alpha\geq 1$, with supermodular binary cost functions, no deterministic online algorithms can achieve
  • $\alpha$-EF1 and completeness,
  • or $\alpha$-MMS and completeness,
  • or $\alpha$-min-USC and completeness.
Proof.
Figure 3: An instance with $\frac{1}{0}$-EF1 allocations
First, let us consider EF1. Consider an online allocation instance; there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1}$ and $c_{2}$.
Based on the relationship between the supermodular and binary functions and the mathematical rank function in Lemma 7, we construct cost functions for the instance by constructing a matroid rank function. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $\{C^{i}_{1},\ldots,C^{i}_{l}\}$. In Figure 3, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 0. W.l.o.g., we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0. We will discuss the following two cases.
Case 1: we allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-1: we allocate $e_{3}$ to agent $1$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. To not break the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{3},e_{6}\}$, $A^{6}_{2}=\{e_{4},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{6}_{1}\setminus\{e\})$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{3}\}$, $A^{6}_{2}=\{e_{4},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$. Case 1-2: we allocate $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-2-1: we allocate $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{4},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{4}\}$, $A^{6}_{2}=\{e_{3},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$. Case 1-2-2: we allocate $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $1$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{5},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{4}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{5}\}$, $A^{6}_{2}=\{e_{3},e_{4},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$.
Case 2: we allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 0. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 2-1: we allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{3},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{4}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{3}\}$, $A^{5}_{2}=\{e_{2},e_{4},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$. Case 2-2: we allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{4},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{3}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{4}\}$, $A^{5}_{2}=\{e_{2},e_{3},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$.
Figure 4: An instance with $\frac{1}{0}$-MMS allocations
Next, let us consider MMS or USC. Consider an online allocation instance; there are two agents $\{1,2\}$ with supermodular binary cost functions $v_{1}v_{2}$. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $C^{i}_{1},\ldots,C^{i}_{l}$. In Figure 4, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 0. We allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0. We will discuss the following two cases.
Case 1: we allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$.
Case 2: we allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 0. Continue to discuss in two sub-cases Case 2-1: we allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. Case 2-2: we allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are 0. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. ∎
4.3 Additive Personalized Bi-valued Costs
In this part, we consider additive personalized bi-valued cost functions. For any agent $i\in N$, her cost function $c_{i}$ is additive personalized bi-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$. Similar to the results of goods setting, even for two agents, an EF1 or MMS allocation cannot be guaranteed.
4.3.1 Two Agents
First, we study the additive personalized bi-valued cost functions for two agents and propose the Adapted Chores Envy-Graph Procedure (Algorithm 9) to return an allocation that achieves the tight approximation guarantee for EF1 and has a good approximation for MMS. Using techniques similar to the case of goods, we take the Chores Envy-Graphs procedure (Algorithm 10) as the base algorithm. When it may not be further executed (an envy cycle will appear), we skip to the Chores Preliminary-Breaking-Cycle algorithm (Algorithm 11), where we carefully allocate the future items and guarantee the desired approximation. During the execution of the Chores Preliminary-Breaking-Cycle (CPBC) algorithm, in one complicated case ($(\lambda_{1},\mu_{1})=(2,2)$), we introduce the Chores Deep-Breaking-Cycle algorithm (Algorithm 12) to address it. The Chores Deep-Breaking-Cycle (CDBC) algorithm may be consecutively executed in many iterations, but the CPBC algorithm is at most consecutively executed in two iterations. Note that the CPBC and CDBC algorithms are used to remove the possible envy cycle and guarantee that in some further iteration, we can return to the CEGP algorithm.
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive personalized bi-valued cost functions
Output: A complete approximate EF1 and approximate MMS allocation $\bf A$}
1
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
Initialize $\alpha=1$; // Algorithm Indicator
3
Initialize $(\lambda_{1},\mu_{1})$ and $(\lambda_{2},\mu_{2})$, where $\lambda_{1}=\lambda_{2}=1$ and $\mu_{1}=\mu_{2}=0$; ;
  // $\lambda$: the times of consecutive execution for the CPBC or CDBC algorithm, and $\mu$: the case label in the CPBC or CDBC algorithm.
4
5Build an envy graph $G$ for two agents;
6 when item $e\in T$ arrives do
7 if $\alpha=1$ then
8
9    $A,\alpha\leftarrow\mathsf{CEGP}(N,A,\alpha,e)$;
10    if $\alpha=2$ then
       $A,\alpha\leftarrow\mathsf{CPBC}(N,A,\alpha,e,(\lambda_{1},\mu_{1}))$; ;
          // Envy cycle will appear.
11
12      end if
13
14 else if $\alpha=2$ then
15    $A,\alpha,(\lambda_{1},\mu_{1})\leftarrow\mathsf{CPBC}(N,A,\alpha,e,(\lambda_{1},\mu_{1}))$;
16 else
17    $A,\alpha,(\lambda_{2},\mu_{2})\leftarrow\mathsf{CDBC}(N,A,\alpha,e,(\lambda_{2},\mu_{2}))$;
18   end if
19
20  Update the envy graph $G$;
21
return $\bf A$};
Algorithm 9 (ACEGP) Adapted Chores Envy-Graph Procedure
Input: $N,{\bf A},\alpha,e$
Output: A complete allocation $\bf A$} and $\alpha$
1
2if  there is no envy between the two agents  then
3 if $\exists i\in N$, $c_{i}(e)=a_{k}$  then
4      Choosing the agent $i$ who values it at $a_{i}$ (breaking ties arbitrarily);
5 else
6      Arbitrarily choose one agent $i$;
7   end if
8
9else
10   Let $i\in N$ be an agent who does not envy the other agent and j=2ij=2-i;
11 if $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ then
12    $\alpha=2$;
13    return $\bf A$} and $\alpha$;
14   end if
15
16 end if
[…]
return $\bf A$} and $\alpha$;
Algorithm 10 (CEGP) Chores Envy-Graph Procedure
Input: $N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1})$
Output: A complete allocation $\bf A$}, $\alpha$ and $(\lambda_{1},\mu_{1})$
1
2if $\lambda_{1}=1$ then
3
4 if $|A_{j}|=1$ then
5
6    $A_{j}=A_{j}\cup\{e\}$;
7    $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
8 else
9
10    if $c_{i}(e)=b_{i}$ then
11
12       $A_{j}=A_{j}\cup\{e\}$;
13       $\lambda_{1}=2$ and $\mu_{2}=1$;
14    else
15       $A_{i}=A_{i}\cup\{e\}$;
16       $\lambda_{1}=2$ and $\mu_{2}=2$;
17      end if
18
19   end if
20
21else
22
23 if $\mu_{2}=1$  then
24
25    $A_{i}=A_{i}\cup\{e\}$;
26 else
27
28    if $c_{j}(e)=a_{j}$ then
29
30       $A_{j}=A_{j}\cup\{e\}$;
31    else
32
33       $A_{i}=A_{i}\cup\{e\}$;
34       $\alpha=3$, $\lambda_{1}=1$, and $\mu_{2}=0$;
35      end if
36
37   end if
38
39 $\alpha=1$, $\lambda_{1}=1$, and $\mu_{2}=0$;
40 end if
41
return $\bf A$}, $\alpha$ and $(\lambda_{1},\mu_{1})$;
Algorithm 11 (CPBC) Chores Preliminary-Breaking-Cycle
Input: $N,{\bf A},\alpha,e,(\lambda_{2},\mu_{2})$
Output: A complete allocation $\bf A$}, $\alpha$, and $(\lambda_{2},\mu_{2})$
1
2if $\lambda_{2}\mod 2=1$ then
3
4 if $c_{j}(e)=a_{j}$ then
5    $A_{j}=A_{j}\cup\{e\}$;
6    $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
7 else if $c_{j}(e)=b_{j}$ and $c_{i}(e)=b_{i}$ then
8
9    $A_{j}=A_{j}\cup\{e\}$;
10    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=1$;
11 else
12    $A_{i}=A_{i}\cup\{e\}$;
13    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=2$;
14   end if
15
16 end if
17else if $\lambda_{2}\mod 2=0$ then
18 if  $\mu_{2}=1$  then
19    if $c_{j}(e)=b_{j}$ then
20
21       $A_{i}=A_{i}\cup\{e\}$;
22       $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
23    else
24       $A_{j}=A_{j}\cup\{e\}$;
25       $\alpha=1$, $\lambda_{2}=1$ and $\mu_{2}=0$;
26      end if
27
28 else
29
30    $A_{j}=A_{j}\cup\{e\}$;
31    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
32   end if
33
34
35return $\bf A$}, $\alpha$, and $(\lambda_{2},\mu_{2})$;
Algorithm 12 (CDBC) Chores Deep-Breaking-Cycle
Definition 11 (Envy Graph for Chores).
Given an allocation ${\bf A}^{k}$, the corresponding envy graph is defined as $G=(V,E)$, where the vertex set $V$ corresponds to agent set $N$, and a directed edge $(i,j)\in E$ iff agent $i$ envies agent $j$, i.e., $c_{i}(A_{i}^{k})>c_{i}(A_{j}^{k})$. Additionally, the directed cycle in the envy graph is called an envy cycle.
Theorem 10.
For the deterministic allocation of indivisible chores for two agents with additive personalized bi-valued cost functions, the Adapted Chores Envy-Graph Procedures (Algorithm 9) computes a complete allocation that satisfies $2$-EF1 and $\frac{5}{3}$-MMS.
Before we show the above Theorem, we first list some lemmas used in the proof.
Lemma 8.
The CEGP algorithm (Algorithm 10) computes an EF1 allocation without an envy cycle.
Proof.
Note that there are three kinds of input allocation. The first one is the initial empty allocation, which is EF1 without an envy cycle. The second one is the output allocation of the CPBC algorithm. The last one is the output allocation of the CDBC algorithm. It suffices to show that the last two kinds of input allocation are EF1 without an envy cycle. That is because if the input allocation of the CEGP algorithm is EF1 without an envy cycle, by the description of the algorithm and induction, it is easy to check that the property of EF1 without an envy cycle is always maintained. By Lemmas 9 and 10, we can derive that when the CPBC or CDBC algorithm outputs $\alpha=1$, implying that the CEGP algorithm is ready to be executed, the output allocation of the CPBC or CDBC algorithm is EF1 without an envy cycle. Therefore, our claim holds. ∎
Observation 4.
When the CPBC algorithm (Algorithm 11) is executed, for the first arriving item $e$, we have $c_{j}(e)=a_{j}$.
Proof.
When the CPBC algorithm is ready to be executed, it means that if the arriving item $e$ is allocated to agent $i$, there will be an envy cycle, i.e, $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ and $c_{i}(A_{i}\cup\{e\})>c_{i}(A_{j})$ For agent $j$, her bundle is EF1. Suppose that $c_{j}(e)=b_{j}$. Then, we have $c_{j}(A_{j})\leq c_{j}(A_{i})+c_{j}(f)\leq c_{j}(A_{i})+b_{j}=c_{j}(A_{i}\cup\{e\})$ for some $f\in A_{j}$, which is a contradiction. Therefore, it must hold that $v_{j}(e)=a_{j}$. ∎
Observation 5.
When the CPBC algorithm (Algorithm 11) is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
Proof.
Suppose that $|A_{i}|=0$. Then, we have $|A_{j}|\geq 1$ since agent $j$ envies agent $i$. After allocating item $e$ to agent $i$, they envies each other, i.e, $c_{i}(e)>c_{i}(A_{j})$ and $c_{j}(A_{j})>c_{j}(e)$. In that case, we derive that $c_{j}(A_{j})=b_{j}$ and $c_{i}(A_{j})=a_{i}$, which contradicts the description of our algorithm. Suppose that $|A_{j}|=0$. Then, we have $|A_{i}|=0$. Otherwise, agent $j$ will not envy agent $i$, after allocating item $e$ to agent $i$. In that case, the envy cycle will not appear. Thus, it hold that $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$. ∎
Lemma 9.
The CPBC algorithm (Algorithm 11) computes a complete allocation satisfying:
  • 2-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
Proof.
The input allocation $\textbf{A}=(A_{1},A_{2})$ of the CPBC algorithm is EF1. According to the description of our algorithm, it is executed in at most two rounds. Then, let us consider different rounds.
$\lambda_{1}=1$. When the first item $e_{1}$ arrives, we have $c_{j}(e_{1})=a_{j}$ by Observation 4. When $|A_{j}|=1$, we have $c_{j}(A_{j})=b_{j}$ since she envies agent $i$. Then, this item is allocated to agent $j$. For agent $i$, her bundle is EF since she does not envy agent $j$. For agent $j$, her new bundle is EF1 since $|A_{j}|=1$ and $c_{j}(e_{1})=a_{j}$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be implemented. When $|A_{j}|>1$, let us consider two subcases about the cost of $e_{1}$.
  • If $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\lambda_{2}=2$ and $\mu_{1}=1$. For agent $j$, her new bundle is $2$-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(A_{i})\geq c_{j}(e_{1})=a_{j}$. For agent $i$, her bundle is EF since she does not envy agent $j$. Thus, this allocation is $2$-EF1.
  • If $c_{j}(e_{1})=a_{i}$, agent $i$ picks it and we set $\lambda_{2}=2$ and $\mu_{2}=2$. For agent $i$, her new bundle is EF1 since after removing $e_{1}$ from her bundle, we have $c_{j}(A_{j})\leq c_{j}(A_{i})$. For agent $j$, her bundle is EF1. Thus, this allocation is EF1 with an envy cycle.
In the above cases, the CPBC algorithm continues to be executed.
$\lambda_{1}=2$. When the second item $e_{2}$ arrives, let us consider two cases about $\mu_{1}$.
  • If $\mu_{1}=1$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $c_{i}(A_{i}\cup\{e_{2}\})\leq c_{i}(A_{i})+b_{i}\leq c_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her bundle is EF1 since we have $v_{j}(A_{j}\setminus\{f\})\leq v_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be executed.
  • If $\mu_{1}=2$, let us consider two subcases about the cost of $e_{2}$. When $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{2})=a_{j}\leq c_{j}(e_{1})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be executed. When $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her bundle is EF1 since she does not pick any item in these two rounds. Thus, this allocation is 2-EF1, and we set $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that we need to implement the CDBC algorithm.
Combining the above cases together, it is not hard to see that the claim holds. ∎
Observation 6.
When the CDBC algorithm (Algorithm 12) is executed, there are two properties about allocation A:
  • $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$;
  • Let $e_{1}$ and $e_{2}$ denote the two new allocated items in the CPBC algorithm, we have $c_{i}(A_{i}\setminus\{e\}\leq 2c_{i}(A_{j})$, where $e\in\arg\max_{f\in\{e_{1},e_{2}\}}c_{i}(f)$.
Proof.
When the CDBC algorithm is executed, the CPBC algorithm is terminated for $\lambda_{1}=\mu_{1}=2$. Then, by the proof of Lemma 9, it is easy to check that statements are correct. ∎
Lemma 10.
The CDBC algorithm (Algorithm 12) computes a complete allocation satisfying:
  • 2-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
Proof.
When the CDBC algorithm is executed, by Observation 6, the input allocation $A=(A_{1},A_{2})$ is 2-EF1, where $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$ and $c_{i}(A_{i}\setminus\{e\})\leq 2c_{i}(A_{j})$ for some $e\in A_{i}$. Let $f_{1}$ and $f_{2}$ denote the two items that are allocated to agent $j$ in the CPBC algorithm and we have $\min\{c_{j}(e_{1}),c_{j}(e_{2})\}=a_{j}$.
$\lambda_{2}=1$. When the item $e_{1}$ arrives at $\lambda_{2}=1$, let us consider three cases of cost functions.
  • When $c_{j}(e_{1})=a_{j}$, agent $j$ picks it and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$. Thus, this allocation is EF1 without an envy cycle, implying that the CEGP algorithm is ready to be executed.
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it. Then, we set $\lambda_{2}=1$ and $\mu_{2}=1$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=b_{j}\leq c_{j}(f_{2})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=a_{j}$, agent $i$ picks it. Then, we set $\lambda_{2}=2$ and $\mu_{2}=2$. For agent $i$, her new bundle is still 2-EF1 since we have $\frac{c_{i}(A_{i}\setminus\{f\}\cup\{e_{1}\})}{c_{i}(A_{j})}\leq\frac{c_{i}(A_{j})+c_{i}(e_{1})}{c_{i}(A_{j})}\leq 2$ for some $f\in\{f_{1},f_{2}\}$ where the inequality follows from the fact that $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}$ and $|A_{j}|\geq 2$. Thus, this allocation is 2-EF1.
It can be seen that in the last two cases, the CDBC algorithm continues to be executed.
$\lambda_{2}=2$. When the item $e_{2}$ arrives at $\lambda_{2}=2$, let us consider two cases about $\mu_{2}$.
  • When $\mu_{2}=1$, we need to consider two subcases. If $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$ and $c_{j}(e_{2})\leq c_{i}(e_{1})=b_{i}$. For agent $j$, her bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{1})=b_{j}=c_{j}(e_{2})$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda_{2}=1$ and $\mu_{2}=0$, implying that the CDBC algorithm continues to be executed. If $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{2})\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=b_{j}=c_{j}(f_{2})$ and $c_{j}(e_{2})=a_{j}=c_{j}(f_{1})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$, implying that the CEGP algorithm is ready to be executed.
  • When $\mu_{2}=1$, agent $j$ picks it, For agent $i$, her bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{2})\leq c_{j}(e_{1})=b_{j}$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda_{2}=1$ and $\mu_{2}=0$, implying that the CDBC algorithm continues to be executed.
Combining the above cases, it can be seen that our statement holds. ∎
Proof of Theorem 10.
In the main body of the CAGEP algorithm 9, it can be seen that for different values of the algorithm indicator $\alpha$, we need to choose the corresponding algorithm to execute. By Lemmas 8, 9, and 10, it can be concluded that the output allocation is complete and $2$-EF1. Regarding MMS, by the implication between EF1 and MMS Sun et al. (2021), i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{n\alpha+n-1}{n-1+\alpha}$-MMS, where $n$ is the number of agents, it is easy to see that the output allocation is $\frac{5}{3}$-MMS. ∎
Next, we complement our approximation results by providing the lower bounds below.
Theorem 11.
For two agents with additive personalized bi-valued cost functions, no deterministic online algorithm can compute a complete allocation that is $(2-\epsilon)$-EF1 or $(\frac{3}{2}-\epsilon)$-MMS for any $\epsilon>0$.
Proof.
$e_{1}$ $e_{2}$ $e_{3}$
agent 11 11 11 $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 1
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 11 11 $\frac{1}{\epsilon}$ 11
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 2.2
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
agent 11 11 11 $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 2.3
Table 10: The impossibility results of EF1 for additive personalized bi-valued cost functions for two agents.
$e_{1}$ $e_{2}$
agent 11 11 11
agent 2 11 11
Case 1
$e_{1}$ $e_{2}$ $e_{3}$
agent 11 11 11 2
agent 2 11 11 22
Case 2
Table 11: The impossibility results of MMS for additive personalized bi-valued cost functions for two agents.
Let $a_{1}=a_{2}=1$ and $b_{1}=b_{2}=\frac{1}{\epsilon}$ where $0<\epsilon<1$. Consider the following online instance for $n=2$, where the cost of each item is shown in Table 12. To guarantee the completeness of the chores allocation, we should guarantee that each arriving item is allocated to one agent.
EF1.
When $t=1$, the arriving item has the cost of 1 for both agents. Without loss of generality, assume that agent 1 picks it. When $t=2$, the item that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. To avoid the approximation ratio of EF1 being infinity, it is allocated to agent 2. Then, it is easy to check that the allocation is EF1 and MMS. In round $t=3$, the item that has the cost of $\frac{1}{\epsilon}$ for both agents arrives. Then, let us consider two cases.
Case 1. If it is allocated to agent 2, no further item will appear. Thus, the allocation is $\frac{1}{\epsilon}$-EF1.
Case 2. If it is allocated to agent 1, it is clear that this allocation is EF1. Next, when $t=4$, the arriving item has the cost of $1$ for agent 1 and $\frac{1}{\epsilon}$ for agent 2. Then, let us consider two subcases. If agent 1 picks it, no further item will appear. So this allocation is trvially $(1+\frac{1}{\epsilon})$-EF1. If agent $2$ picks it, the allocation is EF1. At last, in round $t=5$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. If it is allocated to agent 1, the allocation is $(\frac{1}{2\epsilon}+\frac{1}{2})$-EF1. If it is allocated to agent 2, the allocation is $\frac{2}{1+\epsilon}$-EF1.
Combining the above cases together, we can derive that the lower bound of EF1 is 2.
MMS.
The constructed instance is below. Let $a_{1}=a_{2}=1$, $b_{1}=b_{2}=2$. When $t=1$, the arriving item has the cost of 1. Assume that it is allocated to agent 1. Then, in round $t=2$, the item that has the cost of 1 for both agents arrives. If it is allocated to agent 1, no further allocation will appear. Thus, this allocation is 2-MMS. If it is allocated to agent 2, this allocation is MMS. In round $t=3$, the arriving item has the cost of 2 for both agents. No matter which agent picks it, the final allocation is $\frac{3}{2}$-MMS. Combining the above cases together, we can derive that the lower bound of MMS is $\frac{3}{2}$. ∎
At last, we examine the compatibility between approximate EF1 or MMS and USC, and provide the following negative results.
Proposition 5.
Given any $\alpha,\gamma\geq 1$ and $\beta<2$, for the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation satisfying
  • $\alpha$-EF1 and $\gamma$-min-USC,
  • or $\beta$-MMS and $\gamma$-min-USC,
even for two agents with additive personalized bi-valued cost functions.
Proof.
Without loss of generality, we assume that $c_{1}\in\{1,2\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+\epsilon}{\epsilon+1}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha\geq 1$ or $\beta$-MMS for any $\beta<2$, and $\gamma$-min-USC for any given $\gamma\geq 1$ when two agents have the additive personalized bi-valued cost functions. ∎
4.3.2 The Combination with Additive Binary Costs
Next, we consider the setting where one agent (agent $n$) has the additive personalized bi-valued cost function, and the remaining agents have additive binary cost functions and propose the Adapted-Chores-Picking Algorithm (Algorithm 13) to compute a complete allocation that satisfies EF1 and MMS. The intuition of this algorithm is similar to the goods setting. In Table 12, we illustrate how our algorithm works for three agents. Regarding efficiency, we show that EF1 or MMS is incompatible with USC in this setting.
Input: An instance $(T,N,(c_{i})_{i\in N})$ with (n1)(n-1) additive binary cost functions $(c_{1}.\ldots,c_{n-1})$ and one additive personalized bi-valued cost function $(c_{n})$.
Output: A complete EF1 and MMS allocation $\bf A$}
1
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5  Let $i_{min}\in\arg\min_{i\in[n-1]}|\{f~|f\in A_{i}~and~c_{i}(f)>0\}|$;
6 if $\exists i^{*}\in N$ such that $c_{i^{*}}(e)=0$ then
7
8    $A_{i^{*}}=A_{i^{*}}\cup\{e\}$;
9 else
10
11    if $c_{n}(e)=a_{n}$ then
12
13       if $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\geq|A_{n}|$ then
14          $A_{n}=A_{n}\cup\{e\}$;
15       else
16
17          $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
18         end if
19
20    else
21       if $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\leq|A_{n}|$ then
22
23          $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
24       else
25          $A_{n}=A_{n}\cup\{e\}$;
26         end if
27
28      end if
29
30   end if
31
32
return $\bf A$};
Algorithm 13 Adapted-Chores-Picking Algorithm
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $\ldots$
agent 11 11 0 11 1 11 $\ldots$
agent 22 11 1 1 11 1 $\ldots$
agent 3 $a_{3}$ $a_{3}$ $a_{3}$ $b_{3}$ $b_{3}$ $\ldots$
Table 12: A simple illustration for three agents of Algorithm 13.
Theorem 12.
For the deterministic allocation of indivisible chores for (n1)(n-1) agents with additive binary cost functions and one agent with the additive personalized bi-valued cost function, the Adapted-Chores-Picking Algorithm (Algorithm 13) computes a complete allocation that satisfies EF1 and MMS.
Proof.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 13.
EF1.
We show that the computed allocation is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after the round $k$, the allocation $A^{k}$ is EF1. We show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is also EF1.
  • If there exists an agent $i\in N$ such that $c_{i}(e_{k+1})=0$, agent $i$ picks it. It is clear that this allocation is still EF1.
  • If $c_{n}(e_{k+1})=a_{n}$, then let us consider two subcases. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\geq|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is still EF1. For agent $n$, her bundle is EF1 since she has the priority of picking the item with the cost of $a_{n}$, and the number of items in her bundle is no more than that of items with non-zero costs in agent $i_{min}$’s bundle. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|<|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the additive binary cost valuation and the lowest number of items with non-zero costs.
  • If $v_{n}(e_{k+1})=b_{n}$, the let us consider two subcases. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\leq|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the additive binary cost valuation and the lowest number of items with non-zero costs. For any agent $i\in N\setminus\{i_{min}\}$, her bundle remains untouched and is still EF1. When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|>|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is trivially EF1. For agent $n$, her bundle is EF since she has the priority of picking the item with the cost of $a_{n}$, and in this case, the extra item in another agent’s bundle must have the cost of $b_{n}$ for agent $n$.
Combining the above cases, it can be seen that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k+1}$ is still EF1.
MMS.
Fix an arbitrary agent $i\in N$ and an arbitrary round $k$. If her bundle is EF, we are done. If not, then let us consider two cases. When $i\in[n-1]$, she has the additive binary valuation function, and by our algorithm description, the number of items with non-zero costs in her bundle is at most one more than that of the other agent’s bundle. Thus, it is not hard to see that her bundle is MMS. When $i=n$, currently, the number of items in her bundle is one more than that of the other agent’s bundle, and the extra item has the cost of $a_{n}$. After removing this item, the cost of her bundle is no more than that of the other agent’s bundle from her perspective, so we can conclude that her bundle is MMS. ∎
Proposition 6.
Given any $\alpha,\beta\geq 1$, for the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation satisfying
  • $\alpha$-EF1 and $\beta$-min-USC,
  • or MMS and $\beta$-min-USC,
even for two agents where one agent has the additive binary cost function and the other agent has the additive personalized bi-valued cost function.
Proof.
Without loss of generality, we assume that $c_{1}\in\{0,1\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+1}{2}=\frac{1}{2\epsilon^{2}}+\frac{1}{2}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm outputs a complete allocation that is $\alpha$-EF1 or MMS in the setting of additive binary and additive personalized bi-valued cost functions, and guarantees $\beta$-min-USC for any given $\alpha,\beta\geq 1$. ∎
5 Conclusion and Future Work
In our paper, we study the fair online allocation of indivisible items and present a series of positive and negative results regarding the existence and approximation results in fairness and efficiency with various valuation or cost functions, which provide a comprehensive view of what can be achieved. For future work, the most interesting direction is to fill the gap in the approximation of EF1 and MMS for bi-valued valuation or cost functions when there are more than two agents, because it involves the complicated envy cycle, and the current technique cannot be directly generalized. Additionally, based on our results in other online settings, it would also be interesting to study the items with deadline setting in a more complicated valuation or cost function setting.
Appendix A Other Online Settings
Online algorithms face two roadblocks: unknown future information and the inability to revoke previous decisions. A natural question is how to avoid the worst-case scenario in online algorithms. In this section, we study two special settings. One is that each arriving item can wait for one period and does not need to be allocated immediately, which can be found in the online matching literature Ashlagi et al. [2018]. In this setting, we prove that an EF1 allocation for two agents with additive personalized bi-valued valuation/cost functions exists. The other one is knowing some information in advance, which originates from the online MMS allocation with normalized valuations Zhou et al. [2023]. In this setting, we show that the Marginal-Greedy Algorithm computes a non-wasteful EF1 allocation when items are goods, and the Round-Robin algorithm computes a complete EF1 allocation when items are chores.
A.1 Items with Deadline
In this part, each item has a deadline $d$ after its arrival. After $d$ units of time, we must decide how to allocate it. For two agents with additive personalized bi-valued valuation functions in the goods setting, we design the Adapted Matching Algorithm (Algorithm 14) to compute an NW allocation that satisfies EF1. For two agents with additive personalized bi-valued cost functions in the chores setting, we design the Adapted Chores Matching Algorithm (Algorithm 15) to compute a complete allocation that satisfies EF1. The common intuition of the above two algorithms is that we find a perfect matching between two agents and two items after two consecutive items arrive so that we can guarantee that there is no envy cycle between two agents and EF1 always holds.
Theorem 13.
For the deterministic allocation of indivisible goods for two agents with additive personalized bi-valued valuation functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Matching Algorithm (Algorithm 14) computes an NW allocation that satisfies EF1.
Lemma 11.
In Algorithm 14, for the allocation $A$ after allocating two consecutive online items $e_{1}$ and $e_{2}$, we have the following two properties for any agent $i\in N$:
  • If $v_{i}(A_{i})<v_{i}(A_{j})$, then $v_{i}(A_{i})-v_{i}(A_{j})=a_{i}-b_{i}$;
  • If $v_{i}(A_{i})>v_{i}(A_{j})$, then $v_{i}(A_{i})-v_{i}(A_{j})\geq b_{i}-a_{i}$.
Proof.
We show the claim by induction. For the base case, the statement trivially holds. For the induction step, after allocating items $e_{k-1}$ and $e_{k}$, the two properties hold. Without loss of generality, let $A$ denote the allocation after allocating items $e_{k-1}$ and $e_{k}$. Then, we show that allocating items $e_{k+1}$ and $e_{k+2}$, our claim also holds. Fix an arbitrary agent $i\in N$. Now, consider two cases about her.
  • $v_{i}(A_{i})\geq v_{i}(A_{j})$. Because one agent picks exactly one item, it is clear that after allocating items $e_{k+1}$ and $e_{k+2}$, the claim holds since $v_{i}(e_{k+1}),v_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
  • $v_{i}(A_{i})<v_{i}(A_{j})$. In this case, according to our algorithm description, she picks one item with a value that is greater than or equal to that of the other one. Suppose that she picks $e_{k+1}$. If $v_{i}(e_{k+1})=v_{i}(e_{k+2})$, then we have $v_{i}(A_{i}\cup\{e_{k+1}\})-v_{i}(A_{j}\cup\{e_{k+2}\})=a_{i}-b_{i}+v_{i}(e_{k+1})-v_{i}(e_{k+2})=a_{i}-b_{i}$. If $v_{i}(e_{k+1})>v_{i}(e_{k+2})$, then we have $v_{i}(A_{i}\cup\{e_{k+1}\})-v_{i}(A_{j}\cup\{e_{k+2}\})=a_{i}-b_{i}+v_{i}(e_{k+1})-v_{i}(e_{k+2})=a_{i}-b_{i}+b_{i}-a_{i}=0$.
This completes the induction and establishes the correctness of our claim. ∎
Proof of Theorem 13.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$is the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 14.
EF1.
We will show that the returned allocation A is always EF1 by induction. For the base case, the empty bundle is trivially EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $A^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i. If she picks $e_{k+1}$, we are done. If not, agent $j$ picks it and we get $v_{i}(e_{k+1})=a_{i}$. For agent $j$, she still does not envy agent $i$, and her new bundle is EF. For agent $i$, by Lemma 11, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k})=a_{i}-b_{i}$. Then, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where $A_{i}^{k}$ is EF1 and there must exists an item $g\in A_{j}$ such that $v_{i}(g)=b_{i}$, implying the allocation $A^{k+1}$ is EF1 without an envy cycle. When no envy exists between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i. If agent $i$ picks $e_{k+1}$, we are done. If not, for agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})-v_{i}(A_{j}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where there must exist $g\in A_{j}$ such that $v_{i}(g)=b_{i}$. For agent $j$, she does not envy agent $i$ and by Lemma 11, we have $v_{j}(A_{j})-v_{j}(A_{i})\geq b_{j}-a_{j}$. Thus, her bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i. If she picks it, we are done. If not, it means that item $e_{k}$ is allocated to agent $i$ and $v_{i}(e_{k})\geq v_{i}(e_{k+1})$. For agent $i$, considering that $A_{i}^{k-1}$ is EF1, we can derive that her bundle is still EF1. For agent $j$, it is trivial that her new bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
Combining the above case, we can conclude that the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
Theorem 14.
For the deterministic allocation of indivisible chores for two agents with additive personalized bi-valued cost functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Chores Matching algorithm (Algorithm 15) computes a complete allocation that satisfies EF1.
Lemma 12.
In Algorithm 15, for the allocation A after allocating two consecutive online items $e_{1}$ and $e_{2}$, we have the following two properties for any agent $i\in N$:
  • If $c_{i}(A_{i})<c_{i}(A_{j})$, then $c_{i}(A_{i})-c_{i}(A_{j})\leq a_{i}-b_{i}$;
  • If $c_{i}(A_{i})>c_{i}(A_{j})$, then $c_{i}(A_{i})-c_{i}(A_{j})=b_{i}-a_{i}$.
Proof.
We show the claim by induction. For the base case, the statement trivially holds. For the induction step, after allocating items $e_{k-1}$ and $e_{k}$, the two properties hold. Let A denote such an allocation. Then, we show that after allocating items $e_{k+1}$ and $e_{k+2}$, our claim also holds. Fix an arbitrary agent $i\in N$. Now, consider two cases about her.
  • $c_{i}(A_{i})<c_{i}(A_{j})$. After allocating two items $e_{k+1}$ and $e_{k+2}$, the two properties hold since agent $i$ only picks one item and has a additive personalized bi-valued cost function i.e., $c_{i}(e_{k+1}),c_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
  • $c_{i}(A_{i})>c_{i}(A_{j})$. By our algorithm description, she picks one item with a cost that is lower than or equal to the other one. Suppose that she picks $e_{k+1}$. If $c_{i}(e_{1})=c_{i}(e_{k+2})$, we have $c_{i}(A_{i}\cup\{e_{k+1}\})-c_{i}(A_{j}\cup\{e_{k+2}\})=b_{i}-a_{i}$. If $c_{i}(e_{1})<c_{i}(e_{2})$, we have $c_{i}(A_{i}\cup\{e_{k+1}\})-c_{i}(A_{j}\cup\{e_{k+2}\})=b_{i}-a_{i}+a_{i}-b_{i}=0$.
This completes the induction and establishes the correctness of our claim. ∎
Proof of Theorem 14.
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 15.
EF1.
We will show that the output allocation A is always EF1 by induction. For the base cases, the empty bundle is EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $\textbf{A}^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i. If she picks it, we are done. If agent $j$ picks it, it means that $c_{j}(e_{k+1})=a_{j}$. For agent $i$, she does not envy agent $j$, so her bundle is still EF. For agent $j$, by Lemma 12, we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$. Then, we have $c_{j}(A_{j}^{k}\setminus\{f\}\cup\{e_{k+1}\})-c_{j}(A_{i})=b_{j}-a_{j}-b_{j}+a_{j}=0$ where $A_{j}^{k}$ is EF1 and there must exist an item $f$ such that $c_{j}(f)=b_{j}$, implying that allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. When there is no envy between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
  • $\lambda=2$ and item $e_{k+1}$ needs to allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i. If agent $i$ picks it, we are done. If not, agent $j$ picks it. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$ and $c_{j}(e_{k+1})=a_{j}$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
  • $\lambda=2$ and item $e_{k+2}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i. If agent $i$ picks it, we are done. If not, agent $j$ picks it. In that case, item $e_{k}$ is allocated to agent $i$. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, we can still derive that her new bundle is EF1 since by our assumption, we have $A^{k-1}_{j}$ is EF1, and $c_{j}(e_{k+2})\leq c_{j}(e_{k+1})$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
Combining the above cases, we can conclude that the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive personalized bi-valued valuation functions and deadline $d=1$
Output: An NW EF1 allocation A
1
2Let $A=(\emptyset,\ldots,\emptyset)$;
3Initialize $\lambda=1$;
4 when item $e\in T$ arrives do
5
6 if $\lambda=1$  then
7
8     Let this item wait for $d$ period and use $e_{1}$ to denote it.
9    $\lambda=2$;
10 else
11
12     Let $e_{2}$ denote the new arriving item;
13    if  there exists an envy between the two agents  then
14
15        Let $i\in N$ be the agent who envies the other one and j=2ij=2-i;
16       if  $v_{i}(e_{1})=b_{i}$  then
17          $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
18       else if $v_{i}(e_{2})=b_{i}$ then
19          $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
20       else if $v_{j}(e_{1})=b_{j}$ then
21          $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
22       else
23          $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
24         end if
25
26    else
27       if  $v_{i}(e_{1})>v_{i}(e_{2})$  then
28          $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
29       else if  $v_{i}(e_{1})<v_{i}(e_{2})$  then
30          $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
31       else
32          if  $v_{j}(e_{1})>v_{j}(e_{2})$  then
33             $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
34          else
35             $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
36            end if
37
38         end if
39
40      end if
41
42    $\lambda=1$;
43   end if
44
45
46
47if  $\lambda=1$ and no further items appear  then
48 if there is no envy between two agents then
49
50     Allocate item $e_{1}$ to an arbitrarily agent;
51 else
52
53     Let $i\in N$ be the agent who envies the other one and j=2ij=2-i;
54    if $v_{i}(e_{1})=b_{i}$ then
55       $A_{i}=A_{i}\cup\{e_{1}\}$;
56    else
57       $A_{j}=A_{j}\cup\{e_{1}\}$;
58      end if
59
60   end if
61
62
return A;
Algorithm 14 Adapted Matching Algorithm
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive personalized bi-valued cost functions and deadline $d=1$
Output: A complete EF1 allocation A
1
2Let $A=(\emptyset,\ldots,\emptyset)$;
3Initialize $\lambda=1$;
4 when item $e\in T$ arrives do
5
6 if $\lambda=1$  then
7
8     Let this item wait for $d$ period and use $e_{1}$ to denote it.
9    $\lambda=2$;
10 else
11
12     Let $e_{2}$ denote the new arriving item;
13    if  there exists an envy between the two agents  then
14
15        Let $i\in N$ be the agent who does not envy the other one, and j=2ij=2-i;
16       if  $c_{j}(e_{1})=b_{j}$  then
17          $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
18       else if $c_{j}(e_{2})=b_{j}$ then
19          $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
20       else if $c_{i}(e_{1})=b_{i}$ then
21          $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
22       else
23          $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
24         end if
25
26    else
27       if  $c_{j}(e_{1})>c_{j}(e_{2})$  then
28          $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
29       else if  $c_{j}(e_{1})<c_{j}(e_{2})$  then
30          $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
31       else
32          if  $c_{i}(e_{1})>c_{i}(e_{2})$  then
33             $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
34          else
35             $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
36            end if
37
38         end if
39
40      end if
41
42    $\lambda=1$;
43   end if
44
45
46
47if  $\lambda=1$ and no further items appear  then
48 if there is no envy between two agents then
49
50     Allocate item $e_{1}$ to an arbitrarily agent;
51 else
52
53     Let $i\in N$ be the agent who does not envy the other one and j=2ij=2-i;
54    if $c_{j}(e_{1})=b_{j}$ then
55       $A_{i}=A_{i}\cup\{e_{1}\}$;
56    else
57       $A_{j}=A_{j}\cup\{e_{1}\}$;
58      end if
59
60   end if
61
62
return A;
Algorithm 15 Adapted Chores Matching Algorithm
A.2 Partial Information in Advance
In this part, agents’ preferences over items are the same, and the arriving order of items is known. With this information, we can obtain a non-wasteful EF1 allocation by the Marginal-Greedy (Algorithm 2) in the goods setting and a complete EF1 allocation by the Round-Robin algorithm in the chores setting.
Definition 12.
An online goods/chores instance is monotone if for any agent $i\in N$, we have $f_{i}(e_{1})\geq\cdots\geq f_{i}(e_{T})$ or for any agent $i\in N$, we have $f_{i}(e_{1})\leq\ldots\leq f_{i}(e_{T})$, where $f_{i}\in\{v_{i},c_{i}\}$ .
Theorem 15.
For the deterministic allocation of indivisible goods from a monotone instance, the Marginal-Greedy Algorithm (Algorithm 2) computes an NW allocation that is EF1.
Proof.
Without loss of generality, we assume that the picking sequence is $1,\ldots,n$. First, let us consider the case of $v_{i}(e_{1})\leq\ldots\leq v_{i}(e_{T})$ for any $i\in N$. Note that, for any agent $i\in N$, if she values some items at zero, these items will arrive before the items with non-zero values. In this phase, she does not pick anything until the item with non-zero values arrives. Additionally, when these zero items are allocated to other agents, from the perspective of agent $i$, the value of the other agent’s bundle remains unchanged. Thus, without loss of generality, we assume that $v_{i}(e_{1})\neq 0$ for any agent $i\in N$. Fix an arbitrary round $k$ and an arbitrary agent $i\in N$. Let $\textbf{A}^{k}$ denote the allocation after the allocation of $e_{k}$. Suppose that she picks item $e_{k}$. For agent $i$, we have $v_{i}(A^{k}_{i})\geq v_{i}(A^{k}_{j}\setminus\{e\})$, where $e$ is the last item added to agent $j$’s bundle, since for any $j<i$, agent $i$ is the latter one to pick the item and for any $j>i$, $v_{i}(e_{s})\geq v_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and st=ni+js-t=n-i+j. For agent $j\neq i$, we have $v_{j}(A_{j}^{k})\geq v_{j}(A_{i}^{k}\setminus\{e\})$, where $e$ is the last item added to agent $i$’s bundle, since for any $j>i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and st=jis-t=j-i, and $j<i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and st=nj+is-t=n-j+i. Thus, the allocation $A^{k}$ is EF1. As for the case of $v_{i}(e_{1})\geq\ldots\geq v_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it. ∎
Theorem 16.
For the deterministic allocation of indivisible chores from a monotone instance, the Round-Robin Algorithm computes a complete allocation that is EF1.
Proof.
Without loss of generality, we assume that the picking sequence is $1,\ldots,n$. First, let us consider the case of $c_{i}(e_{1})\leq\ldots\leq c_{i}(e_{T})$ for any $i\in N$. Fix an arbitrary round $k$ and an arbitrary agent $i\in N$. Let $\textbf{A}^{k}$ denote the allocation after the allocation of $e_{k}$. Suppose that she picks item $e_{k}$. For agent $i$, we have $c_{i}(A^{k}_{i}\setminus\{e_{k}\})=c_{i}(A^{k-1}_{i})\leq c_{i}(A^{k-1}_{j})=c_{i}(A^{k}_{j})$ since for any $j>i$, agent $i$ has the priority of picking the item and for any $j<i$, $c_{i}(e_{s})\leq c_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and ts=ni+jt-s=n-i+j. For agent $j\neq i$, we have $c_{j}(A_{j}^{k})\leq c_{j}(A_{i}^{k})$ for any $j<i$ agent $j$ has the priority of picking items and $j>i$, $c_{j}(e_{s})\leq c_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and ts=nj+it-s=n-j+i. Therefore, the allocation $\textbf{A}^{k}$ is EF1. As for the case of $c_{i}(e_{1})\geq\ldots\geq c_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it.∎
In the monotone instance, each agent has the same preference over all items, and the arriving order of items follows the nonincreasing/nondecreasing order in terms of value or cost. Then, if we allow the arriving order to be arbitrary and keep the same preference, we have the following proposition.
Proposition 7.
For an online goods/chores instance where agents only have identical preferences, there exists an instance for which no EF1 allocation can be found, even for two agents with additive personalized bi-valued valuation or cost functions.
Proof.
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 11 11 $\epsilon$ 11
agent 2 11 $\epsilon$ $\epsilon$ 11
Case 1
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
agent 11 11 11 $\epsilon$ 11 11
agent 2 11 $\epsilon$ 1 $\epsilon$ 11
Case 2
Table 13: The nonexistence of EF1 for two agents that have the same preference over goods.
$e_{1}$ $e_{2}$ $e_{3}$
agent 11 11 11 $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 1
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
agent 11 11 11 $\frac{1}{\epsilon}$ 1 $\frac{1}{\epsilon}$
agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 2
Table 14: The nonexistence of EF1 for two agents that have the same preference over chores.
First, we consider the online goods instance. Without loss of generality, we assume that $v_{1},v_{2}\in\{\epsilon,1\}$, where $0<\epsilon<1$ The value of each item is shown in Table 13. When $t=1$, the arriving item has the value of $1$ for both agents. Without loss of generality, it is allocated to agent 1. For the second round $t=2$, the item that has the value of $1$ for agent 1 and $\epsilon$ for agent 2. To guarantee EF1, agent 2 picks it. In round $t=3$, the arriving item has the value of $\epsilon$ for both agents. Then, let us consider two cases related to allocating it.
  • If agent 2 picks it, this allocation is still EF1. In the last round $t=4$, the arriving item has the value of $1$ for both agents. In this case, it is not hard to check that no matter which agent picks it, the resulting allocation is not EF1.
  • If agent 1 picks it, this allocation is still EF1. Then, in round $t=4$, the arriving item has the value of $1$ for agent 1 and $\epsilon$ for agent 2. To guarantee that EF1 is not violated, it should be allocated to agent 2. In this last round $t=5$, the item has the value of $1$ for both agents. In this case, it is trivial that as long as one agent picks it, the final allocation is not EF1.
Next, let us consider the online chores instance. We assume that $c_{1},c_{2}\in\{1,\frac{1}{\epsilon}\}$ where $0<\epsilon<1$. The cost of each item is shown in Table 14. When $t=1$, the arriving item has the cost of 1 for both agents, and we assume that agent 1 picks it. In round $t=2$, the item that has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2. To maintain EF1, agent 2 picks it. For the next round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. Then, let us consider two cases related to allocating it.
  • If it is allocated to agent 2, no further items appear. In this case, it is clear that this allocation is not EF1.
  • If it is allocated to agent 1, in the following round $t=4$, the arriving item has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2. It should be allocated to agent 2 to ensure that the property of EF1 still holds. In the last round $t=5$, this arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. In this case, it is easy to verify that no allocation is EF1 when it is allocated to one agent.
Combining the above cases, we can conclude that our statement holds. ∎